### v7

# On the Construction of Relativistic Quantum Wave Equation

Abstract
Using the elementary axioms of special relativity and quantum mechanics we construct a wave equation which generalizes the Schrodinger equation. We also solve the general second order differential equation ($y^{\prime\prime}(x)=V(x)y(x)$).
1 Introduction
Einstein belived that the complete equivalence of mass and energy is given from the equation
$E=mc^{2},$ (1)
where $m$ is the mass of a moving particle and $E$ the total energy ($c$ is the velocity of light in vacum). This formula is not completely right. It has to do with error terms of series representing the mass of a moving particle (see relation (5) below) and restrictions using Newton formula to write
$m=m_{0}\gamma(v)=m_{0}+\frac{1}{c^{2}}\left(\frac{1}{2}m_{0}v^{2}\right)+\ldots$ (2)
Taking the first two terms of (2) the energy of a moved particle $E=mc^{2}$ can divided into $E_{0}=m_{0}c^{2}$ (which $m_{0}$ is the rest mass) and $\frac{1}{2}m_{0}v^{2}$ is the kinetic energy.
In 1927 Klein and Gordon manage to constuct a relativistic quantum mechanical equation for moving particles using the following equation
$E^{2}=p^{2}c^{2}+m_{0}^{2}c^{4}.$ (3)
Here $p$ is the momentum of a moving particle and $m_{0}$ is the rest mass. By attaching where $E$ and $p$ the quantum operators $i\hbar\partial_{t}$ and $i\hbar\partial_{x}$ respectively (energy and momentum), they arived to an equation which bear their names (Klein-Gordon equation):
$\partial^{2}_{xx}Y-\frac{1}{c^{2}}\partial^{2}_{tt}Y-\frac{m_{0}^{2}c^{2}}{\hbar^{2}}Y=0.$ (4)
In some cases the above equation holds good. But in general is not satisfactory.
However the folowing formula for the mass of a traveling particle is reliable
$m=m_{0}\gamma(v)=\frac{m_{0}}{\sqrt{1-\left(\frac{v}{c}\right)^{2}}}$ (5)
In this article, under the same axioms used by Klein and Gordon and not using energy cuts and avoiding the square in energy we arive to a generalized equation similar to that of Schrodinger. Below we state the two axioms we need:
Axiom 1.
Relation (5) holds for every moving particle.
Axiom 2.
i) In the energy $E$ corespond the quantum operator $E=i\hbar\partial_{t}$
ii) In the momentum $p$ corspond the quantum operator $p=-i\hbar\partial_{x}$.
2 Main Results
The form of the equation we looking for must read as
$EY=\left(f\left(p\right)+E_{0}I\right)Y,$ (6)
where $f$ is a single valued function. $E_{0}=m_{0}c^{2}$, is the energy of the particle and the quantities $m_{0},c$ are the rest mass and the velocity of light in vacum respectively.
We shall work with simple single-valued calculus and operators, to arive to an equation formed by (5) (Axiom 1), then we will use the quantum corespodence (Axiom 2) to produce the desired diferential equation.
Starting right away we have:
$E=\frac{p^{2}}{2m}+m_{0}c^{2}=\frac{m^{2}v^{2}}{2m}+m_{0}c^{2}=\frac{1}{2}mv^{2}+m_{0}c^{2}=\frac{1}{2}m_{0}\gamma(v)v^{2}+m_{0}c^{2}=$
$\frac{m_{0}c^{2}}{2\sqrt{1-\left(\frac{v}{c}\right)^{2}}}\left(\frac{v}{c}\right)^{2}+m_{0}c^{2}.$ (7)
But
$\frac{v}{c}=\frac{p}{mc}=\frac{p}{m_{0}c}\sqrt{1-\left(\frac{v}{c}\right)^{2}}$ (8)
We interested to expand $E$ in a form of $E=f(m_{0},p)$ and then use Axiom (2). As someone can see this is quite dificult because of the nature of the relation $v=p/m$ in which $m$ depends by $v$ (relation (5)) and $v$ by $p/m$ again. In this point we mention that we can solve (8) with respect to $\frac{v}{c}$ to get
$\xi=\frac{v}{c}=\frac{p}{m_{0}c}\sqrt{1-\xi^{2}}.$ (9)
Hence
$\xi=\frac{v}{c}=\pm\frac{p}{m_{0}c\sqrt{1+\frac{p^{2}}{m_{0}^{2}c^{2}}}}.$ (10)
Seting this in (7) we get
$E=\frac{m_{0}c^{2}}{2}\frac{\xi^{2}}{\sqrt{1-\xi^{2}}}+m_{0}c^{2}=\frac{E_{0}}{2}\frac{p^{2}}{m_{0}^{2}c^{2}\sqrt{1+\frac{p^{2}}{m_{0}^{2}c^{2}}}}+E_{0}.$ (11)
Hence from the expansion
$\frac{x^{2}}{\sqrt{1+x^{2}}}=\sum^{\infty}_{n=0}\small\left(\begin{array}[]{cc}-1/2\\ n\end{array}\right)\normalsize x^{2n+2}$
we get
$E=E_{0}+\frac{E_{0}}{2}\sum^{\infty}_{n=0}\small\left(\begin{array}[]{cc}-1/2\\ n\end{array}\right)\normalsize\frac{p^{2n+2}}{(m_{0}c)^{2n+2}}.$
Replacing the arithmetical quntities of $p$ and $E$ with operators (Axiom 2), we get
Theorem 1.
$i\hbar\frac{\partial Y}{\partial t}=E_{0}Y+\frac{E_{0}}{2}\sum^{\infty}_{n=0}(-1)^{n+1}\small\left(\begin{array}[]{cc}-1/2\\ n\end{array}\right)\normalsize\left(\frac{\hbar}{m_{0}c}\right)^{2n+2}\frac{\partial^{2n+2}Y}{\partial x^{2n+2}}.$ (12)
Provited that for all $(x,t)\in\textbf{R}\times\textbf{R}_{+}$, we have
$\lim_{n\rightarrow\infty}\left|\frac{\partial^{2n+2}_{x}Y(x,t)}{\partial^{2n}_{x}Y(x,t)}\right|=l^{2}\in\textbf{R}$ (12.1)
and $l<1/l_{0}$, $l_{0}=\frac{\hbar}{m_{0}c}$.
Proof.
Observe that (from the ratio test)
$r=l_{0}^{2}\frac{2n+1}{2n+2}\left|\frac{\partial^{2n+4}_{x}Y(x,t)}{\partial^{2n+2}_{x}Y(x,t)}\right|<1$
For solving (12), we use the folowing Lemma’s found in [3]:
Lemma 1.
If
$f(x)=\sum^{\infty}_{n=1}A_{n}e^{-nx},$
then
$f(x)=\sum^{\infty}_{n=1}\frac{1}{e^{nx}-1}\sum_{d|n}A_{d}\mu(n/d)$
and the oposite.
Lemma 2.
If $x>0$
$f(x)=\sum^{\infty}_{n=1}\frac{1}{e^{nx}-1}\sum_{d|n}A_{d}\mu(n/d),$
then
$f^{(\nu)}(x)=\sum^{\infty}_{n=1}\frac{1}{e^{nx}-1}\sum_{d|n}A_{d}(-d)^{\nu}\mu(n/d).$
Assume now that a solution have the form
$Y(x,t)=\sum^{\infty}_{k=1}\frac{\sum_{d|k}\chi_{d<1/l_{0}}A_{d}(t)\mu(k/d)}{e^{kx}-1},$ (12.2)
where $l_{0}<1$. Then
$i\hbar\sum^{\infty}_{k=1}\frac{1}{e^{kx}-1}\sum_{d|k}\chi_{d<1/l_{0}}A^{\prime}_{d}(t)\mu(k/d)=E_{0}\sum^{\infty}_{k=1}\frac{\sum_{d|k}\chi_{d<1/l_{0}}A_{d}(t)\mu(k/d)}{e^{kx}-1}+$
$\frac{E_{0}}{2}\sum^{\infty}_{n=0}(-1)^{n+1}C_{-1/2,n}l_{0}^{2n+2}\sum^{\infty}_{k=1}\frac{1}{e^{kx}-1}\sum_{d|k}\chi_{d<1/l_{0}}A_{d}(t)d^{2n+2}\mu(k/d).$ (13)
Rearaging the double series in the above equation and summing with respect to $n$ we get
$i\hbar\sum^{\infty}_{k=1}\frac{1}{e^{kx}-1}\sum_{d|k}\chi_{d<1/l_{0}}A^{\prime}_{d}(t)\mu(k/d)=$
$E_{0}\sum^{\infty}_{k=1}\frac{1}{e^{kx}-1}\sum_{d|k}\chi_{d<1/l_{0}}A_{d}(t)\mu(k/d)-$
$\frac{E_{0}}{2}\sum^{\infty}_{k=1}\frac{1}{e^{kx}-1}\sum_{d|k}\chi_{d<1/l_{0}}A_{d}(t)\mu(k/d)\frac{l_{0}^{2}d^{2}}{\sqrt{1-l_{0}^{2}d^{2}}}.$ (14)
Hence it must be
$\frac{i\hbar}{m_{0}c^{2}}A_{d}^{\prime}(t)=\left(1-\frac{l_{0}^{2}d^{2}}{2\sqrt{1-l_{0}^{2}d^{2}}}\right)A_{d}(t)\textrm{, }d=1,2,\ldots,1/l_{0}$
or equivalently
$A_{n}(t)=f_{n}\exp\left(-\frac{iE_{0}t}{\hbar}+\frac{iE_{0}l_{0}^{2}n^{2}t}{2\hbar\sqrt{1-l_{0}^{2}n^{2}}}\right)\textrm{, for }n=1,2,\ldots,1/l_{0}.$
Hence
$Y(x,t)=\sum^{\infty}_{n=1}\frac{1}{e^{nx}-1}\sum_{d|n}f_{d}\chi_{d<1/l_{0}}\exp\left(-\frac{iE_{0}t}{\hbar}+\frac{iE_{0}l_{0}^{2}d^{2}t}{2\hbar\sqrt{1-l_{0}^{2}d^{2}}}\right)\mu(n/d).$
Returning to Taylor series, we arive to
Theorem 2.
$Y(x,t)=\exp\left(-\frac{iE_{0}t}{\hbar}\right)\sum^{1/l_{0}}_{n=1}f_{n}\exp\left(\frac{iE_{0}l_{0}^{2}n^{2}t}{2\hbar\sqrt{1-l_{0}^{2}n^{2}}}\right)e^{-nx},$ (15)
where it is $Y(x,0)=\sum^{1/l_{0}}_{n=1}f_{n}e^{-nx}=f(x)$ and $l_{0}=\frac{\hbar}{m_{0}c}<1$ is the Compton’s wavelength of the particle.
Notes.
i)
The Compton-Plank lenght $l_{0}$ along with (12.1), has the following meaning: If $l_{0}\geq 1$, then we can not have a solution of the form (12.2), as also the series becomes no convergent and equation (12) is no loger valid. For particles such as the neutrino’s we have Compton length $l_{0}=\frac{\hbar}{m_{\nu}c}\approx 0.000164\textrm{ cm }$. Electron have $l_{0}=3,86\times 10^{-13}\textrm{ cm }$. This in literature is called Compton wavelength of electron.
ii) Taking the first term of the sum in (12) we arive easily to Schrodinger’s equation
$i\hbar\frac{\partial Y}{\partial t}=E_{0}Y-\frac{\hbar^{2}}{2m_{0}}\frac{\partial^{2}Y}{\partial x^{2}}.$ (16)
Set now
$C_{m,n}:=\small\left(\begin{array}[]{cc}m\\ n\end{array}\right)\normalsize.$ (17)
$Y(t;x,y,z)=\sum^{\infty}_{k,l,m=1}\frac{\sum_{d_{1}|k}\sum_{d_{2}|l}\sum_{d_{3}|m}A_{d_{1}d_{2}d_{3}}(t)\mu\left(k/d_{1}\right)\mu(l/d_{2})\mu(m/d_{3})}{(e^{kx}-1)(e^{ly}-1)(e^{mz}-1)}.$
Hence
$\partial^{2\nu+2}_{x}Y(t;x,y,z)=$
$\sum^{\infty}_{k,l,m=1}\frac{\sum_{d_{1}|k}\sum_{d_{2}|l}\sum_{d_{3}|m}A_{d_{1}d_{2}d_{3}}(t)d_{1}^{2\nu+2}\mu\left(k/d_{1}\right)\mu(l/d_{2})\mu(m/d_{3})}{(e^{kx}-1)(e^{ly}-1)(e^{mz}-1)},$
$\partial^{2\kappa+2}_{y}Y(t;x,y,z)=$
$\sum^{\infty}_{k,l,m=1}\frac{\sum_{d_{1}|k}\sum_{d_{2}|l}\sum_{d_{3}|m}A_{d_{1}d_{2}d_{3}}(t)d_{2}^{2\kappa+2}\mu\left(k/d_{1}\right)\mu(l/d_{2})\mu(m/d_{3})}{(e^{kx}-1)(e^{ly}-1)(e^{mz}-1)},$
$\partial^{2\rho+2}_{z}Y(t;x,y,z)=$
$\sum^{\infty}_{k,l,m=1}\frac{\sum_{d_{1}|k}\sum_{d_{2}|l}\sum_{d_{3}|m}A_{d_{1}d_{2}d_{3}}(t)d_{3}^{2\rho+2}\mu\left(k/d_{1}\right)\mu(l/d_{2})\mu(m/d_{3})}{(e^{kx}-1)(e^{ly}-1)(e^{mz}-1)}.$
Hence as in the one dimensional case we have
$i\hbar\sum^{\infty}_{k,l,m=1}\frac{1}{(e^{kx}-1)(e^{ly}-1)(e^{mz}-1)}\sum_{d_{1}|k,d_{2}|l,d_{3}|m}A_{d_{1}d_{2}d_{3}}^{\prime}(t)\mu(k/d_{1})\mu(l/d_{2})\mu(m/d_{3})=$
$E_{0}\sum^{\infty}_{k,l,m=1}\frac{1}{(e^{kx}-1)(e^{ly}-1)(e^{mz}-1)}\sum_{d_{1}|k,d_{2}|l,d_{3}|m}A_{d_{1}d_{2}d_{3}}(t)\mu(k/d_{1})\mu(l/d_{2})\mu(m/d_{3})\times$
$\times\left(1-\frac{l_{0}^{2}d_{1}^{2}}{2\sqrt{1-l_{0}^{2}d_{1}^{2}}}-\frac{l_{0}^{2}d_{2}^{2}}{2\sqrt{1-l_{0}^{2}d_{2}^{2}}}-\frac{l_{0}^{2}d_{3}^{2}}{2\sqrt{1-l_{0}^{2}d_{3}^{2}}}\right).$
Hence it must be
$\frac{i\hbar}{E_{0}}A_{d_{1}d_{2}d_{3}}^{\prime}(t)=\left(1-\frac{l_{0}^{2}d_{1}^{2}}{2\sqrt{1-l_{0}^{2}d_{1}^{2}}}-\frac{l_{0}^{2}d_{2}^{2}}{2\sqrt{1-l_{0}^{2}d_{2}^{2}}}-\frac{l_{0}^{2}d_{3}^{2}}{2\sqrt{1-l_{0}^{2}d_{3}^{2}}}\right)A_{d_{1}d_{2}d_{3}}(t),$
or
$A_{d_{1}d_{2}d_{3}}(t)=$
$f_{d_{1}d_{2}d_{3}}\exp\left[\left(-i\frac{E_{0}t}{\hbar}+i\frac{E_{0}l_{0}^{2}d_{1}^{2}t}{2\hbar\sqrt{1-l_{0}^{2}d_{1}^{2}}}+i\frac{E_{0}l_{0}^{2}d_{2}^{2}t}{2\hbar\sqrt{1-l_{0}^{2}d_{2}^{2}}}+i\frac{E_{0}l_{0}^{2}d_{3}^{2}t}{2\hbar\sqrt{1-l_{0}^{2}d_{3}^{2}}}\right)\right].$
Hence the equation
$i\hbar\frac{\partial Y}{\partial t}=E_{0}Y+\frac{E_{0}}{2}\sum^{\infty}_{n=0}(-1)^{n+1}C_{-1/2,n}\left(\frac{\hbar}{m_{0}c}\right)^{2n+2}\left(\frac{\partial^{2n+2}Y}{\partial x^{2n+2}}+\frac{\partial^{2n+2}Y}{\partial y^{2n+2}}+\frac{\partial^{2n+2}Y}{\partial z^{2n+2}}\right),$
admits solution
$Y(t;x,y,z)=\exp\left(-\frac{iE_{0}t}{\hbar}\right)\times$
$\times\sum^{1/l_{0}}_{k,l,m=1}f_{klm}\exp\left[\frac{iE_{0}l_{0}^{2}t}{2\hbar}\left(\frac{k^{2}}{\sqrt{\left|l_{0}^{2}k^{2}-1\right|}}+\frac{l^{2}}{\sqrt{\left|l_{0}^{2}l^{2}-1\right|}}+\frac{m^{2}}{\sqrt{\left|l_{0}^{2}m^{2}-1\right|}}\right)\right]e^{-kx-ly-mz}.$
QED
Theorem 3.
Assume the equation
$i\hbar\frac{\partial Y}{\partial t}=\frac{E_{0}}{2}\sum^{\infty}_{n=0}(-1)^{n+1}C_{-1/2,n}\left(\frac{\hbar}{m_{0}c}\right)^{2n+2}\frac{\partial^{2n+2}Y}{\partial x^{2n+2}}+E_{0}Y+V(x)Y.$ (17.1)
If the potential $V(x)$ is zero, then
$Y(x,t)=\frac{\exp\left(-\frac{iE_{0}t}{\hbar}\right)}{2\pi}\int^{1/l_{0}}_{-1/l_{0}}\widehat{Y}(\gamma,0)\exp\left(-\frac{iE_{0}t}{2\hbar}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)e^{i\gamma x}d\gamma.$ (18)
If the potential $V(x)$ is a real function of $x$ ($2\pi$-periodic) i.e.
$V(x)=\sum^{\infty}_{k=0}a_{k}x^{k}=\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx},$
then
$i\hbar\frac{\partial}{\partial t}\widehat{Y}(\gamma,t)=\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}+D\right)\widehat{Y}(\gamma,t),$ (19)
where
$D(.)=\sum^{\infty}_{k=0}a_{k}i^{k}\left(\frac{\partial}{\partial\gamma}\right)^{k}(.)=V\left(i\frac{\partial}{\partial\gamma}\right)(.).$ (20)
Hence if we assume the inner product
$\left(\Phi,\Psi\right)=\int^{\infty}_{-\infty}\overline{\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma,$ (20.00)
the Hamiltonian is
$H=E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}+V\left(i\frac{\partial}{\partial\gamma}\right)$ (20.01)
and is self adjoint in all Hilbert space $L_{2}(R)$. If $BL(R)$ denote the space of all Band Limited functions with Fourier transform zero outside of an interval $\sigma,\sigma]$, $\sigma>0$, then $BL(R)$ is Hilbert subspace of $L_{2}(R)$ and also dense in $L_{2}(R)$. Hence when the potential $V(x)$ is real and $2\pi$-periodic function (not limited to $2\pi$) we can write if $Y((.),t)\in BL(R)$:
$i\hbar\frac{\partial}{\partial t}\widehat{Y}(\gamma,t)=H\widehat{Y}(\gamma,t)\textrm{, }\gamma\in[-1/l_{0},1/l_{0}]$ (20.02)
and more generaly (20.02) hold in all $\gamma\in R$. The solution of (20.02) is
$\widehat{Y}(\gamma,t)=\exp\left(-\frac{i}{\hbar}tH\right)\widehat{Y}(\gamma,0).$ (20.03)
Proof.
Assuming the equation
$i\hbar\frac{\partial Y}{\partial t}=\frac{E_{0}}{2}\sum^{\infty}_{n=0}(-1)^{n+1}C_{-1/2,n}\left(\frac{\hbar}{m_{0}c}\right)^{2n+2}\frac{\partial^{2n+2}Y}{\partial x^{2n+2}}+E_{0}Y+V(x)Y,$ (21)
where $Y=Y(x,t)$ and $V(x)=\sum^{\infty}_{k=0}a_{k}x^{k}$. Then taking the Fourier transform in both sides of (21), we get
$i\hbar\frac{\partial\widehat{Y}(\gamma,t)}{\partial t}=\frac{E_{0}}{2}\sum^{\infty}_{n=0}(-1)^{n+1}C_{-1/2,n}l_{0}^{2n+2}(i\gamma)^{2n+2}\widehat{Y}(\gamma,t)+E_{0}\widehat{Y}(\gamma,t)+$
$\int^{\infty}_{-\infty}Y(x,t)V(x)e^{-ix\gamma}dx\textrm{, }|\gamma|\leq 1/l_{0}$ (i)
and $\widehat{Y}(\gamma,t)=\int^{+\infty}_{-\infty}Y(x,t)e^{-ix\gamma}dx$. The only way to hold $(i)$ is to have $\widehat{Y}(\gamma,t)=0$, when $\left|\gamma\right|>1/l_{0}$ i.e. $Y(x,t)$ must be a band limited function. Hence we write equivalent
$i\hbar\frac{\partial\widehat{Y}(\gamma,t)}{\partial t}=\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\widehat{Y}(\gamma,t)+\sum^{\infty}_{n=0}a_{n}i^{n}\frac{\partial^{n}\widehat{Y}(\gamma,t)}{\partial\gamma^{n}}.$ (21.1)
Hence if we denote the differential operator
$D(.)=\sum^{\infty}_{n=0}a_{n}i^{n}\left(\frac{\partial}{\partial\gamma}\right)^{n}(.)=V\left(i\frac{\partial}{\partial\gamma}\right)(.),$ (22)
we get (19). Were we have make use of the periodicity of $V(x)$, ($V(x)$ is bounded function of $x$, since it is periodic function in all $R$) and the fact that $Y((.),t)$ is in $L_{2}(R)$:
$\int^{\infty}_{-\infty}Y(x,t)V(x)e^{-ix\gamma}dx=\int^{\infty}_{-\infty}Y(x,t)\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}e^{-ix\gamma}dx=$
$\sum^{\infty}_{n=-\infty}c_{n}\int^{\infty}_{-\infty}Y(x,t)e^{-ix(\gamma+n)}dx=\sum^{\infty}_{n=-\infty}c_{n}\widehat{Y}(\gamma+n,t)=$
$\sum^{\infty}_{n=-\infty}c_{n}\exp\left(n\frac{\partial}{\partial\gamma}\right)\widehat{Y}(\gamma,t)=V\left(i\frac{\partial}{\partial\gamma}\right)\widehat{Y}(\gamma,t).$
For to solve (19), we write $\widehat{Y}(\gamma,t)=A_{n}(\gamma)B_{n}(t)$. Then
$i\hbar A_{n}(\gamma)B_{n}^{\prime}(t)=\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)A_{n}(\gamma)B_{n}(t)+B_{n}(t)DA_{n}(\gamma).$
Hence
$i\hbar\frac{B_{n}^{\prime}(t)}{B_{n}(t)}=E^{*}_{n}\textrm{ and }DA_{n}(\gamma)=\left(E^{*}_{n}-E_{0}-\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)A_{n}(\gamma).$
Hence
$B_{n}(t)=\Lambda_{n}\exp\left(\frac{-iE^{*}_{n}}{\hbar}t\right)$ (22.1)
and
$\sum^{\infty}_{k=0}a_{k}i^{k}\left(\frac{\partial}{\partial\gamma}\right)^{k}A_{n}(\gamma)=\left(E^{*}_{n}-E_{0}-\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)A_{n}(\gamma)\Leftrightarrow$
$V\left(i\frac{\partial}{\partial\gamma}\right)A_{n}(\gamma)=A_{n}(\gamma)g_{n}(\gamma),$ (eq)
where
$g_{n}(\gamma)=E^{*}_{n}-E_{0}-\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}.$ (22.2)
This equation $(eq)$ is hold for general potential $V(x)$. Assuming that
$V(x)=\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx},$
we have
$V\left(i\frac{\partial}{\partial\gamma}\right)A_{n}(\gamma)=\sum^{\infty}_{k=-\infty}c_{k}\exp\left(k\frac{\partial}{\partial\gamma}\right)A_{n}(\gamma)=\sum^{\infty}_{k=-\infty}c_{k}A_{n}(\gamma+k).$
Hence it is interesting to solve the equation
$\sum^{\infty}_{k=-\infty}c_{k}A_{m}(\gamma+k)=g_{m}(\gamma)A_{m}(\gamma)\textrm{, }m\in\textbf{Z}.$ (23)
We solve the equation:
$\sum^{\infty}_{k=-\infty}c_{k}A_{m}(\gamma+k)=g_{m}(\gamma)A_{m}(\gamma).$ (24)
We have
$\sum^{\infty}_{k=-\infty}c_{k}A_{m}(\gamma+k)=\sum^{\infty}_{k=-\infty}\delta_{k,0}g_{m}(\gamma)A_{m}(\gamma+k)\Leftrightarrow$
$\sum^{\infty}_{k=-\infty}A_{m}(\gamma+k)\left(c_{k}-\delta_{k,0}g_{m}(\gamma)\right)=0.$
If
$A_{m}(\gamma)=\sum_{l\in Z^{*}}b_{m,l}(\gamma)$
and
$b_{m,l}(\gamma+k)=b^{*}_{m,l}(\gamma)G_{m}(kl,\gamma),$
we get
$\sum^{\infty}_{k=-\infty}c_{k}A_{m}(\gamma+k)=c_{0}A_{m}(\gamma)+\sum_{k\in Z^{*}}c_{k}A_{m}(\gamma+k)=$
$c_{0}\sum_{l\in Z^{*}}b_{m,l}(\gamma)+\sum_{(k,l)\in Z^{*}\times Z^{*}}c_{k}b_{m,l}\left(\gamma+k\right)=$
$c_{0}G_{m}(0,\gamma)\sum_{l\in Z^{*}}b^{*}_{m,l}(\gamma)+\sum_{(k,l)\in Z^{*}\times Z^{*}}c_{k}b^{*}_{m,l}(\gamma)G(kl,\gamma)=$
$c_{0}G_{m}(0,\gamma)\sum_{l\in Z^{*}}b^{*}_{m,l}(\gamma)+\sum_{n\in Z^{*}}\left(\sum_{kl=n}c_{k}b^{*}_{m,l}(\gamma)\right)G_{m}(n,\gamma).$
Hence
$c_{0}G_{m}(0,\gamma)\sum_{l\in Z^{*}}b^{*}_{m,l}(\gamma)+\sum_{n\in Z^{*}}G_{m}(n,\gamma)\left(\sum_{kl=n}c_{k}b^{*}_{m,l}(\gamma)\right)=$
$\sum_{n\in Z}G_{m}(n,\gamma)\left(\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}c_{k}b^{*}_{m,l}(\gamma)\right)=\sum_{n\in Z}\sum_{l\in Z^{*}}\delta_{n,0}g_{m}(\gamma)b_{m,l}(\gamma+n)\Leftrightarrow$
$\sum^{\infty}_{n=-\infty}G_{m}(n,\gamma)\left(\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}c_{k}b^{*}_{m,l}(\gamma)\right)=\sum_{n\in Z}\sum_{l\in Z^{*}}\delta_{n,0}g_{m}(\gamma)b^{*}_{m,l}(\gamma)G_{m}(nl,\gamma)=$
$\sum^{\infty}_{n=-\infty}G_{m}(n,\gamma)\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}b^{*}_{m,l}(\gamma)\delta_{k,0}g_{m}(\gamma)\Leftrightarrow$
$\sum^{\infty}_{n=-\infty}G_{m}(n,\gamma)\left(\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}c_{k}b^{*}_{m,l}(\gamma)\right)=$
$\sum^{\infty}_{n=-\infty}G_{m}(n,\gamma)\left(\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}b^{*}_{m,l}(\gamma)\delta_{k,0}g_{m}(\gamma)\right)\Leftrightarrow$
$\sum^{\infty}_{n=-\infty}G_{m}(n,\gamma)\left(\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}b^{*}_{m,l}(\gamma)\left(c_{k}-\delta_{k,0}g_{m}(\gamma)\right)\right)=0.$
Hence
$\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}b^{*}_{m,l}(\gamma)\left(c_{k}-g_{m}(\gamma)\delta_{k,0}\right)=0.$ (24.01)
We set
$a_{m,k}(\gamma)=c_{k}-g_{m}(\gamma)\delta_{k,0}.$
Then if $a_{m,k}(\gamma)$ are known, the arithmetic inverse of $a_{m,k}(\gamma)$, is $b^{*}_{m,k}(\gamma)$ in the sense
$\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}b^{*}_{m,l}(\gamma)a_{m,k}(\gamma)=0\Leftrightarrow b^{*}_{m,n}(\gamma)=\{a_{m,n}(\gamma)\}^{(-1)}$ (24.1)
Equivalent we can find $b^{*}_{m,l}(\gamma)$ from the system:
$\sum_{l\in Z^{*}}b^{*}_{m,l}(\gamma)=0$ (co1)
and
$\sum_{\scriptsize\begin{array}[]{cc}abs(d)|n\\ n\neq 0\end{array}\normalsize}b^{*}_{m,d}(\gamma)c_{n/d}=0$ (co2)
Thus
$A_{m}(\gamma)=G_{m}(0,\gamma)\sum^{\infty}_{n=-\infty}\{c_{n}-g_{m}(\gamma)\delta_{n,0}\}^{(-1)}.$ (25)
Then the solution $\widehat{Y}(\gamma,t)$ of (21.1) is
$\widehat{Y}(\gamma,t)=\sum^{\infty}_{m=-\infty}\Lambda_{m}G_{m}(0,\gamma)\exp\left(\frac{-iE_{m}t}{\hbar}\right)\sum^{\infty}_{n=-\infty}\{c_{n}-g_{m}(\gamma)\delta_{n,0}\}^{(-1)}.$ (26)
Now we show that $H$ is a self adjoint operator in all Hilbert space
$\left(\Phi,H\Psi\right)=\int^{\infty}_{-\infty}\overline{\Phi(\gamma,t)}\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}+V\left(i\frac{\partial}{\partial\gamma}\right)\right)\Psi(\gamma,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\Phi(\gamma,t)}\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\Psi(\gamma,t)d\gamma+$
$\int^{\infty}_{-\infty}\overline{\Phi(\gamma,t)}V\left(i\frac{\partial}{\partial\gamma}\right)\Psi(\gamma,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma+$
$\int^{\infty}_{-\infty}\overline{\Phi(\gamma,t)}\sum^{\infty}_{k=-\infty}c_{k}\exp\left(k\frac{\partial}{\partial\gamma}\right)\Psi(\gamma,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma+$
$\sum^{\infty}_{k=-\infty}c_{k}\int^{\infty}_{-\infty}\overline{\Phi(\gamma,t)}\Psi(\gamma+k,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma+$
$\sum^{\infty}_{k=-\infty}c_{k}\int^{\infty}_{-\infty}\overline{\Phi(\gamma-k,t)}\Psi(\gamma,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma+$
$\int^{\infty}_{-\infty}\overline{\sum^{\infty}_{k=-\infty}\overline{c_{k}}\exp\left(-k\frac{\partial}{\partial\gamma}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma+$
$\int^{\infty}_{-\infty}\overline{\sum^{\infty}_{k=-\infty}\overline{c_{-k}}\exp\left(k\frac{\partial}{\partial\gamma}\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma.$
Hence assuming that $\overline{c_{-k}}=c_{k}$, we get
$\left(\Phi,H\Psi\right)=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}+\sum^{\infty}_{k=-\infty}c_{k}\exp\left(k\frac{\partial}{\partial\gamma}\right)\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma=$
$\int^{\infty}_{-\infty}\overline{\left(E_{0}+\frac{E_{0}}{2}\frac{l_{0}^{2}\gamma^{2}}{\sqrt{1+l_{0}^{2}\gamma^{2}}}+V\left(i\frac{\partial}{\partial\gamma}\right)\right)\Phi(\gamma,t)}\Psi(\gamma,t)d\gamma=\left(H\Phi,\Psi\right).$
But we have asumed that $\overline{c_{-k}}=c_{k}$. Hence equivalent
$\sum^{\infty}_{k=-\infty}\overline{c_{-k}}e^{-ikx}=\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx}\Leftrightarrow$
$\overline{\sum^{\infty}_{k=-\infty}c_{-k}e^{ikx}}=\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx}\Leftrightarrow$
$\overline{\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx}}=\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx}\Leftrightarrow\overline{V(x)}=V(x).$
Hence $V(x)$ must be real for $x$ real. QED
In the most general case we have
$i\hbar\frac{\partial Y(x,t)}{\partial t}=E_{0}Y(x,t)+\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}i^{n}\frac{\partial^{n}Y(x,t)}{\partial x^{n}}+V(x)Y(x,t).$ (26.01)
Assuming that the convergence of the sum in (26.01) is uniform and the function $f(x)=\sum^{\infty}_{n=0}f_{n}x^{n}$ is $f:R\rightarrow R$ is real analytic, then
$H=\frac{E_{0}}{2}f\left(i\frac{\partial}{\partial x}\right)+E_{0}I+V(x)I,$ (26.02)
is self adjoint. That is because we know operator $i\frac{\partial}{\partial x}$ is self-adjoint in all $L_{2}(R)$ and thus the operator $f\left(i\frac{\partial}{\partial x}\right)$, is self adjoint in all $L_{2}(R)$ (from spectral theorem). Hence the equation
$i\hbar\frac{\partial}{\partial t}Y(x,t)=HY(x,t),$ (26.03)
is the semigroup equation of our system and the operator $H$ is self adjoint and independed of $t$. We also have
$Y(x,t)=e^{-i/\hbar tH}Y(x,0).$
Conjecture.
For general functions $f(x),g(x)$, we have the following continuation
$\int^{\infty}_{-\infty}f(x)g(x)e^{-ix\gamma}dx=f\left(i\frac{\partial}{\partial\gamma}\right)\widehat{g}(\gamma)=g\left(i\frac{\partial}{\partial\gamma}\right)\widehat{f}(\gamma).$
Taking the Fourier transform in both sides of $(26.01)$ we have (using and the above notes also)
$i\hbar\frac{\partial\widehat{Y}(\gamma,t)}{\partial t}=E_{0}\widehat{Y}(\gamma,t)+\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}i^{n}\int^{\infty}_{-\infty}\frac{\partial^{n}Y(x,t)}{\partial x^{n}}e^{-ix\gamma}dx+$
$\int^{\infty}_{-\infty}V(x)Y(x,t)e^{-ix\gamma}dx\Leftrightarrow$
$i\hbar\frac{\partial\widehat{Y}(\gamma,t)}{\partial t}=E_{0}\widehat{Y}(\gamma,t)+\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n-1}_{j=0}\left[\left(\frac{\partial^{n-1-j}}{\partial x^{n-1-j}}Y(x,t)\right)e^{-ix\gamma}\right]^{x=+\infty}_{x=-\infty}(i\gamma)^{j}+$
$\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}(-\gamma)^{n}\widehat{Y}(\gamma,t)+V\left(i\frac{\partial}{\partial\gamma}\right)\widehat{Y}(\gamma,t)\Leftrightarrow$
$E\widehat{Y}(\gamma,t)=\left(E_{0}+\frac{E_{0}}{2}f(-\gamma)+V\left(i\frac{\partial}{\partial\gamma}\right)\right)\widehat{Y}(\gamma,t)+$
$\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n-1}_{j=0}\left[\frac{\partial^{n-1-j}}{\partial x^{n-1-j}}Y(x,t)e^{-ix\gamma}\right]^{x=+\infty}_{x=-\infty}(i\gamma)^{j}\Leftrightarrow$
$E\widehat{Y}(\gamma,t)=\left(E_{0}+\frac{E_{0}}{2}f(-\gamma)+V\left(i\frac{\partial}{\partial\gamma}\right)\right)\widehat{Y}(\gamma,t)+$
$\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n-1}_{j=0}\left[Y^{(n-1-j,0)}(h,t)e^{-ih\gamma}\right]^{h=+\infty}_{h=-\infty}(i\gamma)^{j}\Leftrightarrow$
$E\widehat{Y}(\gamma,t)=\left(E_{0}+\frac{E_{0}}{2}f(-\gamma)+V\left(i\frac{\partial}{\partial\gamma}\right)\right)\widehat{Y}(\gamma,t)+$
$\frac{E_{0}}{2}\left[e^{-ih\gamma}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n-1}_{j=0}Y^{(j,0)}(h,t)(i\gamma)^{n-1-j}\right]^{h=+\infty}_{h=-\infty}$. (26.04)
However it holds
$\sum^{n-1}_{j=0}Y^{(j,0)}(h,t)(i\gamma)^{(n-1)-j}=(i\gamma)^{-1}\sum^{n}_{j=0}Y^{(j,0)}(h,t)(i\gamma)^{n-j}$-
$(i\gamma)^{-1}Y^{(n,0)}(h,t).$
Hence
$\left[e^{-ih\gamma}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n-1}_{j=0}Y^{(j,0)}(h,t)(i\gamma)^{n-1-j}\right]^{h=+\infty}_{h=-\infty}$=
$(i\gamma)^{-1}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n}_{j=0}\left[Y^{(j,0)}(h,t)e^{ih\gamma}\right]^{h=+\infty}_{h=-\infty}(i\gamma)^{n-j}-$
$(i\gamma)^{-1}\sum^{\infty}_{n=0}f_{n}i^{n}\left[Y^{(n,0)}(h,t)e^{-ih\gamma}\right]^{h=+\infty}_{h=-\infty}$=
$(i\gamma)^{-1}\left[e^{-ih\gamma}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n}_{j=0}\epsilon_{n-j}Y^{(j,0)}(h,t)(i\gamma)^{n-j}\right]^{h=+\infty}_{h=-\infty}$, (26.05)
where $\epsilon_{n}=1$, if $n=1,2,3,\ldots$ and $\epsilon_{0}=0$. Hence
$\left(\sum^{\infty}_{k=0}Y^{(k,0)}(h,t)x^{k}\right)\frac{i\gamma x}{1-i\gamma x}=\sum^{\infty}_{n=0}x^{n}\sum^{n}_{j=0}Y^{(j,0)}(h,t)\epsilon_{n-j}(i\gamma)^{n-j}.$
Hence if we define the $(TY)_{n}(\gamma,t)$ from the relation
$\left[\frac{xe^{-ih\gamma}}{1-i\gamma x}\sum^{\infty}_{k=0}Y^{(k,0)}(h,t)x^{k}\right]^{h=+\infty}_{h=-\infty}=\sum^{\infty}_{n=0}x^{n}(TY)_{n}(\gamma,t),$ (26.06)
then
$\left[e^{-ih\gamma}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n-1}_{j=0}Y^{(j,0)}(h,t)(i\gamma)^{n-1-j}\right]^{h=+\infty}_{h=-\infty}$=
$(i\gamma)^{-1}\left[e^{-ih\gamma}\sum^{\infty}_{n=0}f_{n}i^{n}\sum^{n}_{j=0}\epsilon_{n-j}Y^{(j,0)}(h,t)(i\gamma)^{n-j}\right]^{h=+\infty}_{h=-\infty}$=
$\sum^{\infty}_{n=0}f_{n}i^{n}(TY)_{n}(\gamma,t).$ (26.07)
Note also that
$(TY)_{n}(\gamma,t)=\left[(i\gamma)^{-1}e^{-ih\gamma}\sum^{n}_{j=0}Y^{(n-j,0)}(h,t)\epsilon_{j}(i\gamma)^{j}\right]^{h=+\infty}_{h=-\infty}$. (26.08)
Theorem 3.01
Assume the equation
$i\hbar\frac{\partial Y(x,t)}{\partial t}=E_{0}Y(x,t)+\frac{E_{0}}{2}\sum^{\infty}_{n=0}f_{n}i^{n}\frac{\partial^{n}Y(x,t)}{\partial x^{n}}+V(x)Y(x,t).$ (26.09)
Then taking the Fourier transform in both sides of $(26.09)$, we find
$E\widehat{Y}(\gamma,t)=\left(E_{0}+\frac{E_{0}}{2}f(-\gamma)+V\left(i\frac{\partial}{\partial\gamma}\right)\right)\widehat{Y}(\gamma,t)+\sum^{\infty}_{n=0}f_{n}i^{n}(TY)_{n}(\gamma,t),$ (26.10)
where the $(TY)_{n}(\gamma,t)$ are defined from the relation
$\left[\frac{xe^{-ih\gamma}}{1-i\gamma x}\sum^{\infty}_{k=0}Y^{(k,0)}(h,t)x^{k}\right]^{h=+\infty}_{h=-\infty}=\sum^{\infty}_{n=0}x^{n}(TY)_{n}(\gamma,t)$ (26.1)
and it holds
$(TY)_{n}(\gamma,t)=\left[(i\gamma)^{-1}e^{-ih\gamma}\sum^{n}_{j=0}Y^{(n-j,0)}(h,t)\epsilon_{j}(i\gamma)^{j}\right]^{h=+\infty}_{h=-\infty}$. (26.2)
The symbol $\epsilon_{n}$ is 1 if $n\neq 0$ and 0 otherwise. Also $f(x)=\sum^{\infty}_{n=0}f_{n}x^{n}$.
The Lorentz transformation is
$x_{k}=\sum^{2}_{\nu=1}\Lambda_{k,\nu}x^{\prime}_{\nu},$ :(Λ)
where $x_{1}=x$, $x_{2}=t$, $x^{\prime}_{1}=x^{\prime}$, $x^{\prime}_{2}=t^{\prime}$ and $\Lambda_{1,1}=\gamma_{0}$, $\Lambda_{1,2}=-\gamma_{0}v$, $\Lambda_{2,1}=-\gamma_{0}v/c^{2}$, $\Lambda_{2,2}=\gamma_{0}$, $\gamma_{0}=\left(1-v^{2}/c^{2}\right)^{-1/2}$.
Lemma.
It holds in general (for any operator $U$ and any function $\Psi$):
$i\hbar\left(\partial_{t}U\right)\Psi=\left[E,U\right]\Psi.$
Theorem 3.1
We assume the equation
$i\hbar\frac{\partial}{\partial t}\Phi(x,t)=H\Phi(x,t),$ (eq)
where the ”Hamiltonian” is
$H=f(x)+V\left(i\frac{\partial}{\partial x}\right).$
If $V(x)=\sum_{n}c_{n}e^{-inx}$, is real periodic potential (here limited to $0,2\pi]$ but not have to in generaly), then $H$ is self adjoint in all Hilbert space $L_{2}(R)$ with inner product
$(X,Y)=\int^{\infty}_{-\infty}\overline{X(x,t)}Y(x,t)dx.$
The solution of $(eq)$ is given from the following eigenvalue problem (since $H$ is self adjoint):
$E\psi_{k}(x,t)=H\psi_{k}(x,t)=E_{k}\psi_{k}(x,t).$
Hence
$f(x)\psi_{k}(x,t)+V\left(i\frac{\partial}{\partial x}\right)\psi_{k}(x,t)=E_{k}\psi_{k}(x,t)$
and
$\psi_{k}(x,t)=\exp\left(-iE_{k}t/\hbar\right)\phi_{k}(x).$
Hence
$\Psi(x,t)=\sum^{\infty}_{k=-\infty}C_{k}\psi_{k}(x,t)=\sum^{\infty}_{k=-\infty}C_{k}\exp\left(-\frac{iE_{k}t}{\hbar}\right)\phi_{k}(x).$
Also then
$V\left(i\frac{\partial}{\partial x}\right)\phi_{k}(x)=\sum^{\infty}_{n=-\infty}c_{n}\phi_{k}(x+n)=g_{k}(x)\phi_{k}(x),$
where
$g_{k}(x)=E_{k}-f(x).$
Hence
$V\left(i\frac{\partial}{\partial x}\right)\phi_{k}(x)=\left(E_{k}-f(x)\right)\phi_{k}(x).$
The functions $\phi_{k}(x)$ and $\psi_{k}(x,t)$ are both orthonormal bases:
$\left(\phi_{k},\phi_{l}\right)=\delta_{k,l}\textrm{ and }\left(\psi_{k},\psi_{l}\right)=\delta_{k,l}.$
We define the sequences of functions $b_{k,l}(x)$, $b^{*}_{k,l}(x)$ and $G_{k}(l,x)$, such that
$b_{k,l}(x+n)=b^{*}_{k,l}(x)G_{k}(nl,x)$
and
$\phi_{k}(x)=\sum_{l\in Z^{*}}b_{k,l}(x).$
Also given the arithmetical functions (with respect to $k$) $\alpha_{k}(x)$, we define the arithmetical functions $\{\alpha_{n}(x)\}^{(-1)}$ as
$\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}\alpha_{k}(x)\beta_{l}(x)=0\Leftrightarrow\beta_{k}(x)=\{\alpha_{k}(x)\}^{(-1)}.$
This is equivalent to say
$\sum_{l\in Z^{*}}\beta_{l}(x)=0$
and
$\sum_{0<abs(d)|n}\alpha_{n/d}(x)\beta_{d}(x)=0.$
However in our problem we have
$\sum_{\scriptsize\begin{array}[]{cc}kl=n\\ l\neq 0\end{array}\normalsize}b^{*}_{m,l}(x)\left(c_{k}-g_{m}(x)\delta_{k,0}\right)=0$
or equivalent
$b^{*}_{m,k}(x)=\{c_{k}-g_{m}(x)\delta_{k,0}\}^{(-1)}.$ (co)
With the above definitions the general problem reduces to:
$\sum_{l\in Z^{*}}b^{*}_{m,l}(x)=0$ (co1)
and
$\sum_{\scriptsize\begin{array}[]{cc}0<abs(d)|n\\ n\neq 0\end{array}\normalsize}c_{n/d}b^{*}_{m,d}(x)=0.$ (co2)
Also writing $Y_{k}(x,t)=A_{k}(x)B_{k}(t)$, we get
$Y(x,t)=\sum^{\infty}_{k=-\infty}C_{k}\psi_{k}(x,t)=\sum^{\infty}_{k=-\infty}\Lambda_{k}Y_{k}(x,t)=$
$\sum^{\infty}_{m=-\infty}\Lambda_{m}\exp\left(\frac{-iE_{m}t}{\hbar}\right)G_{m}(0,x)\sum^{\infty}_{n=-\infty}\{c_{n}-g_{m}(x)\delta_{n,0}\}^{(-1)}.$
Theorem 3.2
Assuming the equation $E\Psi=H\Psi:(eq)$, where
$H=f(x)+V\left(i\frac{\partial}{\partial x}\right)$
and $V(x)$ any smooth periodic function, then $H$ is self adoint. Also, if we take the Lorentz transformation of $(eq)$ to be $E^{\prime}\Psi^{\prime}=H^{\prime}\Psi^{\prime}:(eq)^{\prime}$, then $H^{\prime}$ is self adjoint.
Proof.
After making a Lorentz transformation in $(eq)$, we have
$E^{\prime}\Psi^{\prime}=H^{\prime}\Psi^{\prime}\Leftrightarrow\gamma_{0}E\Psi^{\prime}=H^{\prime}\Psi^{\prime}\Leftrightarrow i\hbar\frac{\partial}{\partial t^{\prime}}\Psi^{\prime}=f(x^{\prime})\Psi^{\prime}+V\left(i\frac{\partial}{\partial x^{\prime}}\right)\Psi^{\prime}\Leftrightarrow$
$\gamma_{0}i\hbar\frac{\partial}{\partial t}\Psi^{\prime}=f(x^{\prime})\Psi^{\prime}+V\left(i\gamma_{0}\frac{\partial}{\partial x}\right)\Psi^{\prime}.$
Since we can assume any (arbitrary) function we want for $f(x)$, we may only prove that the operator $V\left(i\gamma_{0}\partial_{x}\right)$, is self adjoint. Assume that $\Phi=\Phi(x,t)$ and $\Psi=\Psi(x,t)$ are any functions of $L_{2}(R)$. Then
$\left(\Phi,V\left(i\gamma_{0}\frac{\partial}{\partial x}\right)\Psi\right)=\int_{R}\overline{\Phi(x,t)}\sum^{\infty}_{n=-\infty}c_{n}\exp\left(\gamma_{0}n\frac{\partial}{\partial x}\right)\Psi(x,t)dx=$
$\int_{R}\overline{\Phi(x,t)}\sum^{\infty}_{n=-\infty}c_{n}\Psi\left(x+\gamma_{0}n\right)dx=\sum^{\infty}_{n=-\infty}c_{n}\int_{R}\overline{\Phi(x,t)}\Psi(x+\gamma_{0}n,t)dx=$
$\sum^{\infty}_{n=-\infty}c_{n}\int_{R}\overline{\Phi\left(x-\gamma_{0}n,t\right)}\Psi(x,t)dx=$
$\int_{R}\overline{\sum^{\infty}_{n=-\infty}\overline{c_{n}}\exp\left(-n\gamma_{0}\frac{\partial}{\partial x}\right)\Phi(x,t)}\Psi(x,t)dx=$
$\int_{R}\overline{\sum^{\infty}_{n=-\infty}c_{-n}\exp\left(-n\gamma_{0}\frac{\partial}{\partial x}\right)\Phi(x,t)}\Psi(x,t)dx=$
$\int_{R}\overline{\sum^{\infty}_{n=-\infty}c_{n}\exp\left(n\gamma_{0}\frac{\partial}{\partial x}\right)\Phi(x,t)}\Psi(x,t)dx=$
$\int_{R}\overline{V\left(i\gamma_{0}\frac{\partial}{\partial x}\right)\Phi(x,t)}\Psi(x,t)dx=\left(V\left(i\gamma_{0}\frac{\partial}{\partial x}\right)\Phi,\Psi\right).$
QED.
We also have the next
Theorem 4.1
Assuming the equation
$y^{\prime\prime}(x)=V(x)y(x),$ (27)
where $V(x)=\sum^{\infty}_{k=-\infty}c_{k}e^{-ikx}$, we can define the functions $f_{l}(\gamma)$, $f^{*}_{l}(\gamma)$, $G(k,\gamma)$, such that:
$f_{l}(\gamma+k)=f^{*}_{l}(\gamma)G(kl,\gamma)$
and
$\widehat{y}(\gamma)=\sum_{l\in Z^{*}}f_{l}(\gamma)=G(0,\gamma)\sum_{l\in Z^{*}}f^{*}_{l}(\gamma),$ (28)
where $f^{*}_{n}(\gamma)=\{c_{n}+\gamma^{2}\delta_{n,0}\}^{(-1)}$ is the arithmetic inverse of $a_{n}(\gamma)=c_{n}+\gamma^{2}\delta_{n,0}$ in the sense
$\sum_{\scriptsize\begin{array}[]{cc}lk=n\\ l\neq 0\end{array}\normalsize}f^{*}_{l}(\gamma)a_{k}(\gamma)=0.$
Proof.
Assume that $V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}$. We have
$\sum^{\infty}_{k=0}F_{k}y^{(k)}(x)=V(x)y(x)\Leftrightarrow$
$\sum^{\infty}_{k=0}F_{k}\int^{\infty}_{-\infty}y^{(k)}(x)e^{-ix\gamma}dx=\int^{\infty}_{-\infty}y(x)\left(\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}\right)e^{-ix\gamma}dx\Leftrightarrow$
$\sum^{\infty}_{k=0}F_{k}\cdot(i\gamma)^{k}\int^{\infty}_{-\infty}y(x)e^{-ix\gamma}dx=\sum^{\infty}_{n=-\infty}c_{n}\int^{\infty}_{-\infty}y(x)e^{-inx}e^{-ix\gamma}dx\Leftrightarrow$
$F(i\gamma)\widehat{y}(\gamma)=\sum^{\infty}_{n=-\infty}c_{n}\exp\left(n\frac{d}{d\gamma}\right)\left(\int^{\infty}_{-\infty}y(x)e^{-ix\gamma}dx\right)\Leftrightarrow$
$F(i\gamma)\widehat{y}(\gamma)=\sum^{\infty}_{n=-\infty}c_{n}\widehat{y}(\gamma+n).$
Equivalent
$F(i\gamma)\widehat{y}(\gamma)=V\left(i\frac{d}{d\gamma}\right)\widehat{y}(\gamma).$ (30.01)
Hence working as above with $F(x)=x^{2}$, we get
$y^{\prime\prime}(x)=V(x)y(x)\Leftrightarrow V\left(i\frac{d}{d\gamma}\right)\widehat{y}(\gamma)=-\gamma^{2}\widehat{y}(\gamma)\Leftrightarrow$
$\gamma^{2}\widehat{y}(\gamma)=\sum^{\infty}_{n=-\infty}c_{n}\widehat{y}(\gamma+n).$
Hence
$\widehat{y}(\gamma)=\sum_{l\in Z^{*}}f_{l}(\gamma)$
and
$\widehat{y}(\gamma)=G(0,\gamma)\sum^{\infty}_{n=-\infty}\{c_{n}+\gamma^{2}\delta_{n,0}\}^{(-1)}.$
QED.
The equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}y^{(n)}(x)=V(x)y(x),$ (30.1)
where $g_{n}=\frac{g^{(n)}(0)}{n!}$, can be written in the form
$\sum^{\infty}_{n=0}g_{n}\gamma^{n}\widehat{y}(\gamma)=V\left(i\frac{d}{d\gamma}\right)\widehat{y}(\gamma).$
Hence
$V\left(i\frac{d}{d\gamma}\right)\widehat{y}(\gamma)=g(\gamma)\widehat{y}(\gamma),$
which have solution
$\sum^{\infty}_{n=-\infty}c_{n}\widehat{y}(\gamma+n)=g(\gamma)\widehat{y}(\gamma).$
Hence one solution is
$\widehat{y}(\gamma)=\frac{1}{g(\gamma)}\sum_{n_{1}}\frac{c_{n_{1}}}{g(\gamma+n_{1})}\sum_{n_{2}}\frac{c_{n_{2}}}{g(\gamma+n_{1}+n_{2})}\sum_{n_{3}}\frac{c_{n_{3}}}{g(\gamma+n_{1}+n_{2}+n_{3})}\ldots.$
Hence holds the next
Theorem 4.2
The equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x),$
where
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-2\pi inx/T},$
is any $T$-periodic function, have solution $y(x)$ with Fourier transform $\widehat{y}(\gamma)$, such that
$\sum^{\infty}_{n=-\infty}c_{n}\widehat{y}\left(\gamma+\frac{2\pi n}{T}\right)=g(\gamma)\widehat{y}(\gamma).$
The function $g$ is $g(z)=\sum^{\infty}_{n=0}g_{n}z^{n}$. Also in case $T=2\pi$ (not limited to), we have
$\sum^{\infty}_{n=0}b_{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x),$
where
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}$
and
$\sum^{\infty}_{n=-\infty}c_{n}\widehat{y}\left(\gamma+n\right)=B(i\gamma)\widehat{y}(\gamma).$
$\widehat{y}(\gamma)=\frac{1}{B(i\gamma)}\sum_{n_{1}}\frac{c_{n_{1}}}{B(i\gamma+in_{1})}\sum_{n_{2}}\frac{c_{n_{2}}}{B(i\gamma+in_{1}+in_{2})}\sum_{n_{3}}\frac{c_{n_{3}}}{B(i\gamma+in_{1}+in_{2}+in_{3})}\ldots,$
where
$B(x)=\sum^{\infty}_{n=0}b_{n}x^{n}.$
The Fourier transform of $y(x)$ is defined as:
$\widehat{y}(\gamma)=\int^{\infty}_{-\infty}y(t)e^{-it\gamma}dt.$
Set $A_{n}(z)$ and $\theta_{l}(z)$ such that
$\theta_{l}(z)=\sum^{\infty}_{n=-\infty}A_{n}(z)\theta_{l}(z+n)$ (eq1)
and
$\sum^{\infty}_{k=-\infty}c_{k}A_{n-k}(z+k)=\delta_{n,0}g(z),$ (eq2)
we have
$\sum^{\infty}_{k=-\infty}c_{k}\theta_{l}(z+k)=\sum^{\infty}_{k=-\infty}c_{k}\sum^{\infty}_{n=-\infty}A_{n}(z+k)\theta_{l}(z+n+k)=$
$\sum^{\infty}_{k=-\infty}c_{k}\sum^{\infty}_{n=-\infty}A_{n-k}(z+k)\theta_{l}(z+n)=$
$\sum^{\infty}_{n=-\infty}\left(\sum^{\infty}_{k=-\infty}c_{k}A_{n-k}(z+k)\right)\theta_{l}(z+n)=$
$\sum^{\infty}_{n=-\infty}\delta_{n,0}g(z)\theta_{l}(z+n)=g(z)\theta_{l}(z).$
Hence we search for functions $\theta_{l}(z)$ and $A_{n}(z)$ such that $(eq1)$ and $(eq2)$ hold. But equation $(eq2)$ can be written as
$\sum^{\infty}_{n=-\infty}\sum^{\infty}_{k=-\infty}c_{k}A_{n-k}(z+k)e^{inx}=g(z)\sum^{\infty}_{n=-\infty}\delta_{n,0}e^{inx}.$ (29)
Hence
$\sum^{\infty}_{n,k=-\infty}c_{k}A_{n}(z+k)e^{inx}e^{ikx}=g(z).$
Taking in both sides the Fourier transform we find
$\sum^{\infty}_{n=-\infty}\sum^{\infty}_{k=-\infty}c_{k}\widehat{A}_{n}(\gamma)e^{i\gamma k}e^{inx}e^{ikx}=\widehat{g}(\gamma)\Leftrightarrow$
$\overline{V}(x+\gamma)\sum^{\infty}_{n=-\infty}\widehat{A}_{n}(\gamma)e^{inx}=\widehat{g}(\gamma)\Leftrightarrow$
$\sum^{\infty}_{n=-\infty}\widehat{A}_{n}(\gamma)e^{inx}=\frac{\widehat{g}(\gamma)}{\overline{V}(x+\gamma)}.$
Hence
$\widehat{A}_{n}(\gamma)=\widehat{g}(\gamma)\frac{1}{2\pi}\int^{2\pi}_{0}\frac{e^{-itn}}{\overline{V}(t+\gamma)}dt\Leftrightarrow$
$A_{n}(x)=\frac{1}{4\pi^{2}}\int^{\infty}_{-\infty}\widehat{g}(\gamma)\left(\int^{2\pi}_{0}\frac{e^{-itn}}{\overline{V}(t+\gamma)}dt\right)e^{ix\gamma}d\gamma.$
Hence knowing $A_{n}(x)$, then from $(eq1)$ we can find $\theta_{l}(z)$ by iteration i.e.
$\theta_{l}(z)=\sum_{n_{1}}A_{n_{1}}(z)\sum_{n_{2}}A_{n_{2}}(z+n_{1})\sum_{n_{3}}A_{n_{3}}(z+n_{1}+n_{2})\ldots$
and we have the next
Theorem 5.
I.
Assume that
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}$ (30)
and $g(x)$ is a given function. Then equation
$V\left(i\frac{d}{dw}\right)\widehat{y}(w)=\sum^{\infty}_{n=-\infty}c_{n}\widehat{y}(w+n)=g(w)\widehat{y}(w),$ (31)
have solution
$\widehat{y}(w)=\sum_{n_{1}}A_{n_{1}}(w)\sum_{n_{2}}A_{n_{2}}(w+n_{1})\sum_{n_{3}}A_{n_{3}}(w+n_{1}+n_{2})\ldots,$ (32)
where $A_{n}(x)$ is such that
$A_{n}(x)=\frac{1}{4\pi^{2}}\int^{\infty}_{-\infty}\widehat{g}(\gamma)\left(\int^{2\pi}_{0}\frac{e^{-itn}}{\overline{V}(t+\gamma)}dt\right)e^{ix\gamma}d\gamma$ (33)
and $\overline{V}(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}$.
II. When
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}$ (34)
and $g(x)=-x^{2}$, then equation
$y^{\prime\prime}(x)=V(x)y(x),$ (35)
have solution $y(x)$ such that its Fourier transform is
$\widehat{y}(w)=\sum_{n_{1}}A_{n_{1}}(w)\sum_{n_{2}}A_{n_{2}}(w+n_{1})\sum_{n_{3}}A_{n_{3}}(w+n_{1}+n_{2})\ldots,$ (36)
where
$A_{n}(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\delta^{\prime\prime}(\gamma)\left(\int^{2\pi}_{0}\frac{e^{-itn}}{\overline{V}(t+\gamma)}dt\right)e^{ix\gamma}d\gamma.$ (37)
Remarks.
1)
The Fourier transform of $f(x)$ is defined as
$\widehat{f}(x)=\int^{\infty}_{-\infty}f(t)e^{-itx}dt\Leftrightarrow f(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\widehat{f}(\gamma)e^{i\gamma x}d\gamma.$ (38)
2) The function $\delta(x)$ is the Dirac delta function, while $\delta_{k,\nu}$ is the Kronecker’s delta symbol, i.e. $\delta_{k,\nu}=1$ if $k=\nu$ and $\delta_{k,\nu}=0$ if $k\neq\nu$.
3) From
$\frac{1}{2\pi}\int^{\infty}_{-\infty}\delta^{\prime\prime}(w)F(w)e^{iwx}dw=\frac{1}{2\pi}\left(-x^{2}F_{n}(0)+2ixF_{n}^{\prime}(0)+F_{n}^{\prime\prime}(0)\right),$
we get
$A_{n}(x)=\frac{1}{2\pi}\int^{2\pi}_{0}\left(-\frac{x^{2}}{\overline{V}(t)}-\frac{2ix\overline{V}^{\prime}(t)}{\overline{V}(t)^{2}}+\frac{2\overline{V}^{\prime}(t)^{2}}{\overline{V}(t)^{3}}-\frac{\overline{V}^{\prime\prime}(t)}{\overline{V}(t)^{2}}\right)e^{-itn}dt.$
Example 1.
Assume $V(x)=1$, then $c_{n}=\delta_{n,0}$. Hence from (39) we get
$A_{n}(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\delta^{\prime\prime}(\gamma)2\pi\delta_{n,0}e^{i\gamma x}d\gamma=-x^{2}\delta_{n,0}.$ (a)
Hence $\theta(z)=-z^{2}\theta(z)$. Hence one leads to $\theta(z)=\delta(z-i)$. Hence $\widehat{y}(z)=\delta(z-i)$, which gives $y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\delta(z-i)e^{-izx}dz=\frac{e^{x}}{2\pi}\textrm{ : }(a)$. Hence indeed one solution of $y^{\prime\prime}(x)=V(x)y(x)$, with $V(x)=1$ is $y(x)=\frac{e^{x}}{2\pi}$. However one observes that working with functions of $L_{2}(\textbf{R})$ is better than working with distributions. In this example the Fourier integral $(a)$ is over the real line and the distribution is $\delta(x-i)$. Hence xix-i is never 0 at the real line. However many books use this Fourier pair (actually we need here a generalization of the Fourier transform or the notion of distribution functions).
Example 2.
Assume $V(x)=e^{-ix}$. Hence $c_{n}=\delta_{n,1}$. In this case $\overline{V}(x)=e^{ix}$ and
$\frac{1}{2\pi}\int^{2\pi}_{0}\frac{e^{-itn}}{e^{i(t+\gamma)}}dt=2\pi e^{-i\gamma}\delta_{n,-1}.$
Hence
$A_{n}(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\delta^{\prime\prime}(\gamma)2\pi e^{-i\gamma}\delta_{n,-1}e^{i\gamma x}d\gamma=-(x-1)^{2}\delta_{n,-1}.$
Hence from
$\sum^{\infty}_{n=-\infty}c_{n}\theta(z+n)=-z^{2}\theta(z)\Leftrightarrow\theta(z+1)=-z^{2}\theta(z),$
we get easily
$\theta(z)=e^{-i\pi z}\Gamma(z)^{2}.$
Therefore, there exists a solution of $y^{\prime\prime}(x)=e^{-ix}y(x)$ such that its Fourier transform is
$\widehat{y}(w)=e^{-i\pi w}\Gamma(w)^{2}.$
However this evaluation is not in $L_{2}(\textbf{R})$.
Remarks.
If we assume that $y(it)=(Mg)(t)$, where $(Mg)(t)$ is the Mellin transform of $g(t)$, then the Fourier transform of a function is
$\widehat{y}(x)=\int^{+\infty}_{-\infty}y(t)e^{-itx}dt=\int^{+\infty/i}_{-\infty/i}y(it)e^{-i(it)x}idt=$
$i\int^{+i\infty}_{-i\infty}y(it)e^{tx}dt=-i2\pi i\frac{1}{2\pi i}\int^{+i\infty}_{-i\infty}Mg(t)\left(e^{-x}\right)^{-t}dt=$
$2\pi g\left(e^{-x}\right).$
Hence $\widehat{y}\left(x\right)=2\pi g\left(e^{-x}\right)$ iff $y\left(ix\right)=(Mg)(x)$.
Example 3.
Assume that
$g(x)=\frac{1}{x^{2}+1},$
then
$(Mg)(s)=\frac{\pi}{2}\csc\left(\frac{\pi s}{2}\right)\textrm{, }0<Re(s)<2.$
Thus
$\widehat{y}(\gamma)=\frac{2\pi}{1+e^{-2\gamma}}$
and
$y(x)=\frac{\pi}{2}\csc\left(\frac{-i\pi x}{2}\right).$
3 About the complete solution of the equation with periodic potentials and some notes on special cases and the equation $y^{\prime\prime}(x)=V(x)y(x)$.
Assume now that
$\frac{1}{V(x)}=\sum^{\infty}_{n=-\infty}c^{*}_{n}e^{-inx}.$ (39)
Then
$\frac{1}{2\pi}\int^{2\pi}_{0}\frac{e^{-itn}}{\overline{V}(t+\gamma)}dt=\frac{1}{2\pi}\int^{2\pi}_{0}\left(\sum^{\infty}_{k=-\infty}c^{*}_{k}e^{-itn}e^{ik(t+\gamma)}\right)dt=$
$\sum^{\infty}_{k=-\infty}c^{*}_{k}\left(\frac{1}{2\pi}\int^{2\pi}_{0}e^{it(k-n)}dt\right)e^{ik\gamma}=\sum^{\infty}_{k=-\infty}c^{*}_{k}\delta_{k,n}e^{ik\gamma}=c^{*}_{n}e^{in\gamma}$
and
$A_{n}(x)=\int^{\infty}_{-\infty}\delta^{\prime\prime}(\gamma)c^{*}_{n}e^{in\gamma}e^{ix\gamma}d\gamma=-(n+x)^{2}c^{*}_{n}.$
Theorem 5.1
If we denote $c_{n}^{*}$ such that
$\frac{1}{V(x)}=\sum^{\infty}_{n=-\infty}c^{*}_{n}e^{-inx},$ (41.1)
then
I. If $g(x)=-x^{2}$, then
$A_{n}(x)=-(n+x)^{2}c^{*}_{n}.$ (41.2)
II. If $g(x)$ is any analytic function, then
$A_{n}(x)=g(n+x)c^{*}_{n}$ (41.3)
and the solution of
$\sum_{n}c_{n}\widehat{y}(z+n)=g(z)\widehat{y}(z),$ (41.4)
is
$\widehat{y}(z)=\lim_{N\rightarrow\infty}\sum_{n_{1}}\sum_{n_{2}}\ldots\sum_{n_{N}}g(n_{1}+z)g(n_{1}+n_{2}+z)\ldots$
$\ldots g(n_{1}+n_{2}+\ldots+n_{N}+z)c^{*}_{n_{1}}c^{*}_{n_{2}}\ldots c^{*}_{n_{N}}.$ (41.5)
Using (38) we get
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\sum_{n_{1}}\sum_{n_{2}}\sum_{n_{3}}\ldots\sum_{n_{N}}(-1)^{N}(n_{1}+w)^{2}(n_{2}+n_{1}+w)^{2}(n_{3}+n_{2}+n_{1}+w)^{2}\ldots$
$\ldots(n_{N}+n_{N-1}+\ldots+n_{2}+n_{1}+w)^{2}c^{*}_{n_{1}}c^{*}_{n_{2}}\ldots c^{*}_{n_{N}}=$
$\lim_{N\rightarrow\infty}\int^{\infty}_{-\infty}\int^{\infty}_{-\infty}\ldots\int^{\infty}_{-\infty}\delta^{\prime\prime}(t_{1})\delta^{\prime\prime}(t_{2})\ldots\delta^{\prime\prime}(t_{N})\times$
$\times\frac{1}{V(t_{1}+t_{2}+\ldots+t_{N})}\frac{1}{V(t_{2}+t_{3}+\ldots+t_{N})}\ldots\frac{1}{V(t_{N-1}+t_{N})}\frac{1}{V(t_{N})}\times$
$\times e^{-it_{1}w}e^{-it_{2}w}\ldots e^{-it_{N}w}dt_{1}dt_{2}\ldots dt_{N}\Rightarrow$
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\int^{\infty}_{-\infty}\int^{\infty}_{-\infty}\ldots\int^{\infty}_{-\infty}\delta^{\prime\prime}(t_{1})\delta^{\prime\prime}(t_{2})\ldots\delta^{\prime\prime}(t_{N})\times$
$\times\frac{1}{V(t_{1})}\frac{1}{V(t_{1}+t_{2})}\ldots\frac{1}{V(t_{1}+t_{2}+\ldots+t_{N})}\times$
$\times e^{-it_{1}w}e^{-it_{2}w}\ldots e^{-it_{N}w}dt_{1}dt_{2}\ldots dt_{N}=$
$\lim_{N\rightarrow\infty}\frac{\partial^{2}}{\partial h_{1}^{2}}\frac{\partial^{2}}{\partial h_{2}^{2}}\ldots\frac{\partial^{2}}{\partial h_{N}^{2}}\left(f_{N}(h_{1},h_{2},\ldots,h_{N})e^{-ih_{1}w-ih_{2}w-\ldots-ih_{N}w}\right)_{h_{1}=h_{2}=\ldots=h_{N}=0},$
where
$f_{N}(t_{1},t_{2},\ldots,t_{N})=\frac{1}{V(t_{1})V(t_{1}+t_{2})\ldots V(t_{1}+t_{2}+\ldots+t_{N})}.$
Also if we denote
$\epsilon_{n_{1},n_{2},\ldots,n_{N}}:=2^{n_{1}mod2+n_{2}mod2+\ldots+n_{N}mod2}.$
and
$f_{N}^{(n_{1},n_{2},\ldots,n_{N})}=\left(\frac{\partial^{n_{1}+n_{2}+\ldots+n_{N}}}{\partial t_{1}^{n_{1}}\partial t_{2}^{n_{2}}\ldots\partial t_{N}^{n_{N}}}f_{N}(t_{1},t_{2},\ldots,t_{N})\right)_{t_{1}=t_{2}=\ldots=t_{N}=0},$
then
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\sum^{2}_{n_{1},n_{2},\ldots,n_{N}=0}(-1)^{n_{1}+n_{2}+\ldots+n_{N}}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(n_{1},n_{2},\ldots,n_{N})}\cdot(iw)^{2N-n_{1}-n_{2}-\ldots-n_{N}}.$
Hence
$\widehat{y}\left(w\right)=\lim_{N\rightarrow\infty}(iw)^{2N}\sum^{2}_{n_{1},n_{2},\ldots,n_{N}=0}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(n_{1},n_{2},\ldots,n_{N})}\cdot(-iw)^{-n_{1}-n_{2}-\ldots-n_{N}}.$
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\sum^{2N}_{k=0}\left(\sum_{\scriptsize\begin{array}[]{cc}0\leq n_{1},n_{2},\ldots,n_{N}\leq 2\\ n_{1}+n_{2}+\ldots+n_{N}=k\end{array}\normalsize}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(n_{1},n_{2},\ldots,n_{N})}\right)(-iw)^{2N-k}.$
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\sum^{2N}_{k=0}\left(\sum_{\scriptsize\begin{array}[]{cc}0\leq n_{1},n_{2},\ldots,n_{N}\leq 2\\ n_{1}+n_{2}+\ldots+n_{N}=2N-k\end{array}\normalsize}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(n_{1},n_{2},\ldots,n_{N})}\right)(-iw)^{k}.$ (40)
But
$\left(\frac{\partial^{n_{1}+n_{2}+\ldots+n_{N}}}{\partial x_{1}^{n_{1}}\partial x_{2}^{n_{2}}\ldots\partial x_{N}^{n_{N}}}f_{N}(P)\right)_{P=(0,0,\ldots,0)}=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}\normalsize}\int_{\textbf{\scriptsize R\normalsize}}\ldots\int_{\textbf{\scriptsize R\normalsize}}\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})\times$
$\times(i\gamma_{1})^{n_{1}}(i\gamma_{2})^{n_{2}}\ldots(i\gamma_{N})^{n_{N}}d\gamma_{1}d\gamma_{2}\ldots d\gamma_{N}.$ (41)
Hence
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\sum^{2N}_{k=0}\left(\sum_{\scriptsize\begin{array}[]{cc}0\leq n_{1},n_{2},\ldots,n_{N}\leq 2\\ n_{1}+n_{2}+\ldots+n_{N}=2N-k\end{array}\normalsize}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(n_{1},n_{2},\ldots,n_{N})}\right)(-iw)^{k}=$
$\lim_{N\rightarrow\infty}\sum^{2N}_{k=0}\left(\sum_{\scriptsize\begin{array}[]{cc}0\leq n_{1},n_{2},\ldots,n_{N}\leq 2\\ n_{1}+n_{2}+\ldots+n_{N}=k\end{array}\normalsize}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(2-n_{1},2-n_{2},\ldots,2-n_{N})}\right)(-iw)^{k}=$
$\lim_{N\rightarrow\infty}\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})\times$
$\times\sum^{2N}_{k=0}\left(\sum_{\scriptsize\begin{array}[]{cc}0\leq n_{1},n_{2},\ldots,n_{N}\leq 2\\ n_{1}+n_{2}+\ldots+n_{N}=2N-k\end{array}\normalsize}\epsilon_{n_{1},n_{2},\ldots,n_{N}}(i\gamma_{1})^{n_{1}}(i\gamma_{2})^{n_{2}}\ldots(i\gamma_{N})^{n_{N}}\right)(-iw)^{2N-k}\times$
$\times d\gamma_{1}d\gamma_{2}\ldots d\gamma_{N}.$
Hence we get the next
Theorem 6.
Assume the equation
$y^{\prime\prime}(x)=V(x)y(x).$ (42)
Then
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\left(\frac{\partial^{2}}{\partial h_{1}^{2}}\frac{\partial^{2}}{\partial h_{2}^{2}}\ldots\frac{\partial^{2}}{\partial h_{N}^{2}}\frac{e^{-ih_{1}w-ih_{2}w-\ldots-ih_{N}w}}{V(h_{1})V(h_{1}+h_{2})\ldots V(h_{1}+h_{2}+\ldots+h_{N})}\right)_{h_{1}=h_{2}=\ldots=h_{N}=0}.$ (43)
Also if we set
$\widehat{y}_{N}(w):=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})(\gamma_{1}-w)^{2}(\gamma_{2}-w)^{2}\ldots(\gamma_{N}-w)^{2}d\gamma_{1}d\gamma_{2}\ldots d\gamma_{N},$ (44)
then
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\widehat{y}_{N}(w)$ (45)
and
$\widehat{y}(w)=\lim_{N\rightarrow\infty}\sum^{2N}_{k=0}\left(\sum_{\scriptsize\begin{array}[]{cc}0\leq n_{1},n_{2},\ldots,n_{N}\leq 2\\ n_{1}+n_{2}+\ldots+n_{N}=k\end{array}\normalsize}\epsilon_{n_{1},n_{2},\ldots,n_{N}}f_{N}^{(2-n_{1},2-n_{2},\ldots,2-n_{N})}\right)(-iw)^{k}.$ (47.1)
However
$\frac{d^{2N}}{dw^{2N}}\widehat{y}_{N}(w)=\textrm{costant}.$
Hence exists costants $C_{N,k}$, such that
$\widehat{y}_{N}(w)=\sum^{2N}_{k=0}C_{N,k}w^{k}.$
Set
$Y(x):=\frac{1}{V(x)}.$
We have
$\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})=\int_{\textbf{\scriptsize R}^{N}}Y(t_{1})Y(t_{1}+t_{2})\ldots Y(t_{1}+t_{2}+\ldots+t_{N})e^{-it_{1}\gamma_{1}-it_{2}\gamma_{2}-it_{N}\gamma_{N}}dt_{1}dt_{2}\ldots dt_{N}.$ (47.2)
Also
$\widehat{Y}(x)=\int_{\textbf{\scriptsize R}}Y(t)e^{-itx}dt.$
Then from (46) we get
$\widehat{y}_{N}(w)=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}-t_{2})\widehat{Y}(t_{2}-t_{3})\ldots\widehat{Y}(t_{N-1}-t_{N})\widehat{Y}(t_{N}+w)t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}.$
Hence
$\widehat{y}_{N}(w)=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}+w)\widehat{Y}(t_{2}-t_{1})\widehat{Y}(t_{3}-t_{2})\ldots\widehat{Y}(t_{N-1}-t_{N-2})\widehat{Y}(t_{N}-t_{N-1})\times$
$\times t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}=$
$\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}+w)\widehat{Y}(t_{2}-t_{1})\ldots\widehat{Y}(t_{N-1}-t_{N-2})\widehat{Y}(t_{N}-t_{N-1})t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}.$
Hence
$\frac{1}{2\pi}\int_{\textbf{\scriptsize R}}\widehat{y}_{N}(w)e^{iww^{\prime}}dw=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}-t_{2})\widehat{Y}(t_{2}-t_{3})\ldots\widehat{Y}(t_{N-1}-t_{N})\times$
$\times\left(\frac{1}{2\pi}\int_{\textbf{\scriptsize R}}\widehat{Y}(t_{N}+w)e^{iww^{\prime}}dw\right)t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}.$
Hence
$y_{N}(w^{\prime})=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}-t_{2})\widehat{Y}(t_{2}-t_{3})\ldots\widehat{Y}(t_{N-1}-t_{N})\times$
$\times\left(\frac{1}{2\pi}\int_{\textbf{\scriptsize R}}\widehat{Y}(w)e^{i(w-t_{N})w^{\prime}}dw\right)t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}.$
Hence
$y_{N}(w^{\prime})=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}-t_{2})\widehat{Y}(t_{2}-t_{3})\ldots\widehat{Y}(t_{N-1}-t_{N})\times$
$\times Y(w^{\prime})e^{-it_{N}w^{\prime}}t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}\Rightarrow$
$y_{N}(w^{\prime})=Y(w^{\prime})\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}-t_{2})\widehat{Y}(t_{2}-t_{3})\ldots\widehat{Y}(t_{N-1}-t_{N})e^{-it_{N}w^{\prime}}t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}.$
Hence
$\frac{y_{N}(w)}{Y(w)}=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{1}-t_{2})\widehat{Y}(t_{2}-t_{3})\ldots\widehat{Y}(t_{N-1}-t_{N})e^{-it_{N}w}t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}.$
Hence
$\frac{y_{N}(w)}{Y(w)}=\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}^{N}}\widehat{Y}(t_{2}-t_{1})\widehat{Y}(t_{3}-t_{2})\ldots\widehat{Y}(t_{N}-t_{N-1})e^{-it_{1}w}t_{1}^{2}t_{2}^{2}\ldots t_{N}^{2}dt_{1}dt_{2}\ldots dt_{N}=$
$\frac{1}{(2\pi)^{N}}\int_{\textbf{\scriptsize R}}\int_{\textbf{\scriptsize R}}\ldots\int_{\textbf{\scriptsize R}}e^{i(t_{2}-t_{1})w}\widehat{Y}(t_{2}-t_{1})t_{1}^{2}dt_{1}e^{i(t_{3}-t_{2})w}\widehat{Y}(t_{3}-t_{2})t_{2}^{2}dt_{2}\ldots\times$
$\times e^{i(t_{N}-t_{N-1})w}\widehat{Y}(t_{N}-t_{N-1})t_{N-1}^{2}dt_{N-1}e^{-it_{N}w}t_{N}^{2}dt_{N}.$
Hence if we define the operator $F_{n}(x,w)$ such that
$F_{n+1}(x,w)=\frac{1}{2\pi}\int_{\textbf{\scriptsize R}}F_{n}(t,w)e^{(x-t)iw}\widehat{Y}(x-t)t^{2}dt,$ (47.2)
then
$\widehat{F}_{N+1}(\gamma,w)=-\frac{1}{2\pi}e^{i\gamma w}\widehat{Y}(\gamma)\frac{d^{2}}{d\gamma^{2}}\widehat{F}_{N}(\gamma,w)$
and
$\frac{y_{N}(w)}{Y(w)}=\left[\int_{\textbf{\scriptsize R}}F_{N}(t,w)e^{-it\gamma}t^{2}dt\right]_{\gamma=w}=-\left[\frac{d^{2}}{d\gamma^{2}}\widehat{F}_{N}(\gamma,w)\right]_{\gamma=w}.$
Hence
$\frac{y_{\infty}(w)}{Y(w)}=\frac{y(w)}{Y(w)}=-\frac{d^{2}}{d\gamma^{2}}\left[-\frac{1}{2\pi}e^{i\gamma w}\widehat{Y}(\gamma)\frac{d^{2}}{d\gamma^{2}}\left[-\frac{1}{2\pi}e^{i\gamma w}\widehat{Y}(\gamma)\frac{d^{2}}{d\gamma^{2}}\left[\ldots\right]\right]\right]_{\gamma=w}$
and we get the next
Theorem 7.
Assume the equation
$y^{\prime\prime}(x)=V(x)y(x),$ (46)
where the potential $V(x)$ is of the form
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}.$ (47)
If
$Y(x):=\frac{1}{V(x)},$ (48)
then it holds
$y(x)=-Y(x)\frac{d^{2}}{d\gamma^{2}}\left[-\frac{1}{2\pi}e^{i\gamma x}\widehat{Y}(\gamma)\frac{d^{2}}{d\gamma^{2}}\left[-\frac{1}{2\pi}e^{i\gamma x}\widehat{Y}(\gamma)\frac{d^{2}}{d\gamma^{2}}\left[\ldots\right]\right]\right]_{\gamma=x}.$ (49)
Also and obviously one can see that immediately from (48):
$y(x)=\frac{1}{V(x)}\frac{d^{2}}{dx^{2}}\left[\frac{1}{V(x)}\frac{d^{2}}{dx^{2}}\left[\frac{1}{V(x)}\frac{d^{2}}{dx^{2}}\left[\frac{1}{V(x)}\ldots\right]\right]\right].$ (50)
It is of very interest to find the Fourier transform $\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})$. From relation (47.2) and discusion below of it, we find (setting $t_{1}=t_{N}^{\prime}$, $t_{2}=t_{N-1}^{\prime}$, $t_{k}=t_{N-k+1}^{\prime}$ and then $t_{1}^{\prime}=h_{1}-h_{2}$, $t_{2}^{\prime}=h_{2}-h_{3}$, $\ldots$, $t_{N-2}^{\prime}=h_{N-2}-h_{N-1}$, $t_{N-1}^{\prime}=h_{N-1}-h_{N}$, $t_{N}^{\prime}=h_{N}$).
$\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})=\int_{\textbf{\scriptsize R\normalsize}^{N}}Y(t_{N}^{\prime})Y(t_{N}^{\prime}+t_{N-1}^{\prime})\ldots\times$
$\times\ldots Y(t_{N}^{\prime}+t_{N-1}^{\prime}+\ldots+t_{2}^{\prime})Y(t_{N}^{\prime}+t_{N-1}^{\prime}+\ldots+t_{2}^{\prime}+t_{1}^{\prime})\times\ldots$
$\times\exp{\left(-it_{N}^{\prime}\gamma_{1}-it_{N-1}^{\prime}\gamma_{2}-\ldots-it_{2}^{\prime}\gamma_{N-1}-it_{1}^{\prime}\gamma_{N}\right)}dt_{1}^{\prime}dt_{2}^{\prime}\ldots dt_{N}^{\prime}=$
$\int_{\textbf{\scriptsize R}^{N}}Y(h_{N})Y(h_{N-1})\ldots Y(h_{2})Y(h_{1})\times$
$\times\exp{\left(-i\gamma_{1}h_{N}-i(h_{N-1}-h_{N})\gamma_{2}-\ldots-i(h_{2}-h_{1})\gamma_{N-1}-i(h_{1}-h_{2})\gamma_{N}\right)}dh_{1}dh_{2}\ldots dh_{N}=$
$\int_{\textbf{\scriptsize R}^{N}}Y(h_{N})Y(h_{N-1})\ldots Y(h_{2})Y(h_{1})\times$
$\times\exp[-ih_{N}(\gamma_{1}-\gamma_{2})-ih_{N-1}(\gamma_{2}-\gamma_{3})$-$\ldots\times$
$\times\ldots-ih_{3}(\gamma_{N-2}-\gamma_{N-1})-ih_{2}(\gamma_{N-1}-\gamma_{N})-ih_{1}\gamma_{N}]dh_{1}dh_{2}\ldots dh_{N}=$
$\widehat{Y}(\gamma_{1}-\gamma_{2})\widehat{Y}(\gamma_{2}-\gamma_{3})\ldots\widehat{Y}(\gamma_{N-2}-\gamma_{N-1})\widehat{Y}(\gamma_{N-1}-\gamma_{N})\widehat{Y}(\gamma_{N}).$
Hence we get the next
Theorem 8.
Assume
$f_{N}(x_{1},x_{2},\ldots,x_{N})=Y(x_{1})Y(x_{1}+x_{2})\ldots Y(x_{1}+x_{2}+\ldots+x_{N}).$ (52.1)
Then
$\widehat{f}_{N}(\gamma_{1},\gamma_{2},\ldots,\gamma_{N})=\widehat{Y}(\gamma_{1}-\gamma_{2})\widehat{Y}(\gamma_{2}-\gamma_{3})\ldots\widehat{Y}(\gamma_{N-2}-\gamma_{N-1})\widehat{Y}(\gamma_{N-1}-\gamma_{N})\widehat{Y}(\gamma_{N}),$ (52.2)
where
$Y(x)=\frac{1}{V(x)}.$
Theorem 8.1
Assume the real $2\pi$-periodic potential
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}\textrm{ and }\frac{1}{V(x)}=\sum^{\infty}_{n=-\infty}c^{*}_{n}e^{-inx}\textrm{, }x\in\textbf{R}.$
If the equation is
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x)$ (eq)
and $y_{l}(x)$ is a solution of $(eq)$ and $\widehat{y}_{l}(z)=\theta_{l}(z)$, then equations $(eq1),(eq2)$ become equivalent with one equation (see Theorem 5.1):
$\theta_{l}(z)=\sum^{\infty}_{n=-\infty}c^{*}_{n}g(z+n)\theta_{l}(z+n).$ (eq0)
Remarks.
If we demand $\theta_{l}(z)$ to be 11-periodic, we have
$\sum^{\infty}_{n=-\infty}c^{*}_{n}g(z+n)=1.$
The oposite states that if $g(z)=\sum_{n=0}^{\infty}g_{n}z^{n}$ is such that
$\sum^{\infty}_{n=-\infty}c^{*}_{n}g(z+n)=1,$
then $\theta_{l}(z)=\widehat{y}_{l}(z)$ and $y_{l}(x)$ satisfies $(eq)$, if $\theta_{l}(z)$ satisfies $(eq0)$. But if $\theta_{l}(z)$ is 11-periodic then this is the case.
Assuming the notation of Theorem 8.1 we have: If
$\theta(z+1)=H(z)\theta(z),$
then
$\theta(z+n)=\left(\prod_{j=0}^{n-1}H(z+j)\right)\theta(z).$
Hence
$\theta(z)=\sum^{\infty}_{n=0}c^{*}_{n}g(z+n)\theta(z+n)\Leftrightarrow\theta(z)=\sum^{\infty}_{n=0}c^{*}_{n}g(z+n)\left(\prod^{n-1}_{j=0}H(z+j)\right)\theta(z).$
Hence
$1=\sum_{n=0}^{\infty}c^{*}_{n}B_{n}(z),$
where
$B_{n}(z)=g(z+n)\prod^{n-1}_{j=0}H(z+j).$
Hence
$g(z+n)=\frac{B_{n}(z)}{\prod^{n-1}_{j=0}H(z+j)}\Leftrightarrow$
$B_{n+1}(z)=g(z+n+1)\prod^{n}_{j=0}H(z+j)=g(n+z+1)H(z+n)\prod^{n-1}_{j=0}H(z+j)=$
$g(z+n+1)H(z+n)\frac{B_{n}(z)}{g(z+n)}.$
Hence
$B_{n+1}(z)=\frac{g(z+n+1)}{g(z+n)}H(z+n)B_{n}(z).$
Also is very easy to show that
Lemma.
$B_{n+1}(z)=B_{n}(z+1)H(z),$
is equivalent to
$B_{n}(z)=g(z+n)\prod^{n-1}_{j=0}H(z+j),$
where $B_{0}(z)=g(z)$ and $\theta(z+1)=H(z)\theta(z)$.
Proof.
We have
$B_{n+1}(z)\theta(z)=B_{n}(z+1)\theta(z+1).$
Hence if we set $\xi_{n}(z):=B_{n}(z)\theta(z)$, then
$\xi_{n+1}(z)=\xi_{n}(z+1).$
Hence
$\xi_{n}(z)=\xi_{n-1}(z+1)\Leftrightarrow\xi_{n}(z)=\xi_{0}(z+n)\Leftrightarrow$
$B_{n}(z)\theta(z)=B_{0}(z+n)\theta(z+n).$
Hence $B_{n}(z)=g(z+n)\prod^{n-1}_{j=0}H(z+j)$.
If the function $\theta(z)$ is defined from the equation
$\theta(z+1)=H(z)\theta(z),$
then $\theta(z)$ satisfies
$\sum^{\infty}_{n=0}c_{n}\theta(z+n)=g(z)\theta(z),$
iff
$\theta(z)=\sum^{\infty}_{n=0}c^{*}_{n}g(z+n)\theta(z+n),$
iff
$\sum^{\infty}_{n=0}B_{n}(z)=1,$ (a)
where
$B_{n}(z)=c_{n}^{*}g(z+n)\prod^{n-1}_{j=0}H(z+j).$
But this last equation is equivalent to:
$B_{n+1}(z)=\frac{c^{*}_{n+1}}{c^{*}_{n}}H(z)B_{n}(z+1).$ (b)
Hence from $(a),(b)$ we get
$H(z)=\sum^{\infty}_{n=0}\frac{c^{*}_{n}}{c^{*}_{n+1}}B_{n+1}(z).$ (c)
However we need to show
$\frac{c^{*}_{n}}{c^{*}_{n+1}}\frac{B_{n+1}(z)}{B_{n}(z+1)}=H(z)=\textrm{independent of }n.$
For this, always exist function $\lambda_{n}(z)$ such that
$\frac{B_{n+1}(z)}{B_{n}(z)}=\lambda_{n}(z)$
and also
$\frac{c_{n}^{*}}{c^{*}_{n+1}}\frac{B_{n+1}(z)}{B_{n}(z+1)}=\frac{c_{n+1}^{*}}{c_{n+2}^{*}}\frac{B_{n+2}(z)}{B_{n+1}(z+1)}\Leftrightarrow\frac{c^{*}_{n+1}B_{n+2}(z)}{c_{n}^{*}B_{n+1}(z)}=\frac{c^{*}_{n+2}B_{n+1}(z+1)}{c^{*}_{n+1}B_{n}(z+1)}\Leftrightarrow$
$\frac{\frac{1}{c^{*}_{n+2}}B_{n+2}(z)}{\frac{1}{c^{*}_{n+1}}B_{n+1}(z)}=\frac{\frac{1}{c^{*}_{n+1}}B_{n+1}(z+1)}{\frac{1}{c^{*}_{n}}B_{n}(z+1)}.$
Hence if we set
$B^{*}_{n}(z)=\frac{1}{c^{*}_{n}}B_{n}(z)$
and
$\lambda^{*}_{n}(z)=\frac{B^{*}_{n+1}(z)}{B^{*}_{n}(z)},$
then it must be
$\lambda^{*}_{n+1}(z)=\lambda^{*}_{n}(z+1)\Leftrightarrow\lambda^{*}_{n}(z)=\lambda^{*}_{n-1}(z+1).$
Hence iterating we get equivalent
$\lambda^{*}_{n}(z)=\lambda^{*}_{0}(z+n)$
and
$\frac{B^{*}_{n+1}(z)}{B^{*}_{n}(z)}=\lambda^{*}_{0}(z+n)\Leftrightarrow B^{*}_{n}(z)=\left(\prod^{n-1}_{j=0}\lambda^{*}_{0}(z+j)\right)B^{*}_{0}(z).$
Hence equivalent
$H(z)=\frac{B^{*}_{n+1}(z)}{B^{*}_{n}(z+1)}=\frac{B^{*}_{0}(z)\prod^{n}_{j=0}\lambda^{*}_{0}(z+j)}{B^{*}_{0}(z+1)\prod^{n-1}_{j=0}\lambda^{*}_{0}(z+1+j)}=\frac{B^{*}_{0}(z)}{B^{*}_{0}(z+1)}\lambda^{*}_{0}(z).$
Hence
$\lambda^{*}_{0}(z)=\frac{B^{*}_{0}(z+1)}{B^{*}_{0}(z)}H(z).$
Hence
$\frac{B^{*}_{n}(z)}{B^{*}_{0}(z)}=\prod^{n-1}_{j=0}\lambda^{*}_{0}(z+j)=$
$\frac{B^{*}_{0}(z+1)}{B^{*}_{0}(z)}H(z)\frac{B^{*}_{0}(z+2)}{B^{*}_{0}(z+1)}H(z+1)\frac{B^{*}_{0}(z+3)}{B^{*}_{0}(z+2)}H(z+2)\times\ldots$
$\ldots\times\frac{B^{*}_{0}(z+n)}{B^{*}_{0}(z+n-1)}H(z+n-1)=\frac{B^{*}_{0}(z+n)}{B^{*}_{0}(z)g(z+n)}g(z+n)\prod^{n-1}_{j=0}H(z+j)=$
$\frac{B^{*}_{0}(z+n)}{B^{*}_{0}(z)g(z+n)}B^{*}_{n}(z).$
Hence it must be
$B^{*}_{0}(z+n)=g(z+n)\textrm{, }\forall z\in\textbf{R},\forall n\in\textbf{N}\Leftrightarrow B^{*}_{0}(z)=g(z).$
For example for any functions $f_{1}(z),f_{2}(z)$ we have (this give us $(a)$):
$B_{n}(z)=\frac{1}{n!}\sum^{n}_{k=0}C_{n,k}\frac{f_{1}^{(k)}(z)f_{2}^{(n-k)}(z)}{f_{1}(z+1)f_{2}(z+1)},$
where
$C_{n,k}:=\left(\begin{array}[]{cc}n\\ k\end{array}\right),$
is the binomial symbol.
Moreover equation $(b)$ is equivalent to equation
$\frac{B^{*}_{n+1}(z)}{B^{*}_{n}(z)}=\lambda^{*}_{0}(z+n),$
where
$H(z)=\frac{g(z)}{g(z+1)}\lambda^{*}_{0}(z)$
and from $(c)$:
$H(z)=\frac{g(z)}{g(z+1)}\lambda^{*}_{0}(z)=\sum^{\infty}_{n=0}c^{*}_{n}B^{*}_{n+1}(z)=\sum^{\infty}_{n=0}\frac{c^{*}_{n}}{c^{*}_{n+1}}B_{n+1}(z).$
$\frac{B^{*}_{n+1}(z)}{B^{*}_{n}(z)}=\lambda^{*}_{n}(z)=\lambda_{0}^{*}(z+n)=\frac{c_{n}^{*}}{c_{n+1}^{*}}\lambda_{n}(z).$
$\frac{B_{n+1}(z)}{B_{n}(z)}=\lambda_{n}(z)=\frac{c_{0}^{*}}{c_{1}^{*}}\frac{c^{*}_{n+1}}{c^{*}_{n}}\lambda_{0}(z+n).$
$\lambda_{0}^{*}(z)=\frac{c_{0}^{*}}{c_{1}^{*}}\lambda_{0}(z).$
$B_{n}(z)=\left(\frac{c_{0}^{*}}{c_{1}^{*}}\right)^{n}c_{n}^{*}g(z)\prod^{n-1}_{j=0}\lambda_{0}(z+j)$
and
$B_{n}(z)=c_{n}^{*}g(z+n)\prod^{n-1}_{j=0}H(z+j).$
$\lambda_{0}^{*}(z)=\frac{g(z+1)}{g(z)}H(z).$
Hence we have the next theorem:
Theorem 8.2
Given the equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x),$ (eq)
with potential $V(x)=\sum_{n}c_{n}e^{-inx}$ a $2\pi$-periodic function and $1/V(x)=\sum_{n}c_{n}^{*}e^{-inx}$ and $g(x)=\sum^{\infty}_{n=0}g_{n}x^{n}$, we assume that exists a family of functions $B_{n}(z)$ such that:
I.
$\frac{B_{n+1}(z)}{B_{n}(z)}=\frac{c^{*}_{0}}{c^{*}_{1}}\frac{c^{*}_{n+1}}{c^{*}_{n}}\lambda_{0}(z+n),$ (eq1)
where $\lambda_{0}(z)$ is a simple function. And
II.
$\sum^{\infty}_{n=0}B_{n}(z)=1\textrm{, }\forall z\in D,$ (eq2)
$B_{0}(z)=g(z).$ (eq3)
Then
$\frac{c_{0}^{*}}{c_{1}^{*}}\lambda_{0}(z)=\frac{g(z+1)}{g(z)}H(z).$ (eq4)
Hence we find $H(z)$ and determine $\theta(z)$, since
$\frac{\theta(z+1)}{\theta(z)}=H(z).$ (eq5)
Hence we get $\widehat{y}(z)=\theta(z)$. Hence we solve $(eq)$.
4 Some interesting solutions
We know that equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x),$
can be written in the form
$\sum^{\infty}_{n=0}c_{n}\widehat{y}(z+n)=g(z)\widehat{y}(z),$
where $g(x)=\sum^{\infty}_{n=0}g_{n}x^{n}$, when
$V(x)=\sum^{\infty}_{n=0}c_{n}e^{-inx}.$
Also if
$Y(x)=\frac{1}{V(x)}=\sum^{\infty}_{n=-\infty}c^{*}_{n}e^{-inx},$
then
$\widehat{y}(z)=\lim_{N\rightarrow\infty}\sum_{n_{1}\geq 0}\sum_{n_{2}\geq 0}\sum_{n_{3}\geq 0}\ldots\times$
$\times\ldots\sum_{n_{N}\geq 0}A_{n_{1}}(z)A_{n_{2}}(z+n_{1})A_{n_{3}}(z+n_{1}+n_{2})\ldots A_{n_{N}}(z+n_{1}+n_{2}+\ldots+n_{N}),$ (52.3)
where
$A_{n}(x)=g(n+x)c_{n}^{*}.$
Also
$g(z)=\sum_{k}c_{k}\prod^{k-1}_{j=0}f_{k}(z+j)$
Hence $\theta(z)$ is solution of (30.1) and consequently
$\widehat{y}(z)=\theta(z),$
where
$\frac{\theta(z+k)}{\theta(z)}=\prod^{k-1}_{j=0}f_{k}(z+j).$
Hence if $f(z)=z$, then we get $\theta(z)=\Gamma(z)$. If $f(z)=z^{2}$, then $\theta(z)=\Gamma(z)^{2}$. If $f(z)=\frac{z^{2}}{z-1}$, then $\theta(z)=(z-1)\Gamma(z)$ . If $f(z)=2z+1$, then $\theta(z)=2^{z}\Gamma(z+\frac{1}{2})$. If $f(z)=2z^{2}+z$, then $\theta(z)=2^{z}\Gamma(z+\frac{1}{2})\Gamma(z)$. If $f(z)=2z^{2}+z+1$, then $\theta(z)=2^{z}\Gamma(z-\rho_{1})\Gamma(z-\rho_{2})$, where $\rho_{1}=\frac{-1-i\sqrt{7}}{4}$, $\rho_{2}=\frac{-1+i\sqrt{7}}{4}$. By this way we get the next
Theorem 9.
Suppose $P(z)=A\prod^{\nu}_{l=1}(z-\rho_{l})$, $Q(z)=B\prod^{\mu}_{l=1}(z-\sigma_{l})$ and $R(z)=\sum^{\infty}_{l=0}a_{l}z^{l}$, are two polynomials of degree $\nu$, $\mu$ and an entire function respectively. $P,Q$, have no common roots and $H(z)=Cz^{m}e^{-R(z)}P(z)/Q(z)$. Then if
$g(z)=\sum^{\infty}_{k=0}c_{k}\left(\frac{A}{B}\right)^{k}\prod^{k-1}_{j=0}H(z+j)=\sum^{\infty}_{k=0}\prod^{k-1}_{j=0}\frac{A}{B}c^{{}^{\prime}}_{j}H(z+j),$ (51)
where $\prod^{k-1}_{j=0}c_{j}^{{}^{\prime}}=c_{k}$, the equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}y^{(n)}(x)=V(x)y(x)\textrm{, }V(x)=\sum^{\infty}_{n=0}c_{n}e^{-inx},$ (52)
can be set in the form
$\sum^{\infty}_{n=0}c_{n}\widehat{y}(z+n)=g(z)\widehat{y}(z)$ (53)
and a solution of (55) is
$\widehat{y}(z)=\left(\frac{A}{B}\right)^{z}C^{z}\Gamma(z)^{m}e^{-R_{1}(z)}\frac{\prod^{\nu}_{j=1}\Gamma(z-\rho_{j})}{\prod^{\mu}_{j=1}\Gamma(z-\sigma_{j})}.$ (54)
If $R(z)$ is a polynomial of degree $\tau$, then the function $R_{1}(z)$ is a polynomial of degree $\tau+1$ (see notes below). If $\tau=\infty$, then again we use the notes below.
Hence the solution of (54) is
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\left(\frac{A}{B}\right)^{z}C^{z}\Gamma(z)^{m}e^{-R_{1}(z)}\left(\frac{\prod^{\nu}_{j=1}\Gamma(z-\rho_{j})}{\prod^{\mu}_{j=1}\Gamma(z-\sigma_{j})}\right)e^{izx}dz.$ (eq)
Notes.
I.
Actualy if $R(z)=z$, then $R_{1}(z)=\frac{z^{2}}{2}-\frac{z}{2}$. If $R(z)=z^{2}$, then $R_{1}(z)=\frac{z^{3}}{3}-\frac{z^{2}}{2}+\frac{z}{6}$. If $R(z)=z^{3}$, then $R_{1}(z)=\frac{z^{4}}{4}-\frac{z^{3}}{2}+\frac{z^{2}}{4}$, $\ldots$ etc. In general if $R(z)=z^{\tau}$ and $p_{\tau}(n)=\sum^{n-1}_{j=0}j^{\tau}=\sum^{\tau+1}_{k=0}h_{k}n^{k}$, then $R(z)=\sum^{\tau+1}_{k=0}h_{k}z^{k}$. Obviously this procedure is linear i.e. if for example $R(z)=z^{2}-2z$, then $R_{1}(z)=\frac{z^{3}}{3}-\frac{z^{2}}{2}+\frac{z}{6}-2\left(\frac{z^{2}}{2}-\frac{z}{2}\right)=\frac{z^{3}}{3}-\frac{3z^{2}}{2}+\frac{7z}{6}$.
II. If $H(z)$ is meromorphic, with $\rho_{j}$, $j=1,2,\ldots$ roots and $\sigma_{j}$, $j=1,2,\ldots$ poles (infinite roots and poles) and exists entire function $R(z)$ such that:
$H(z)=Cz^{m}e^{-R(z)}\prod^{\infty}_{j=1}\left(\frac{z-\rho_{j}}{z-\sigma_{j}}\right),$ (56.00)
provided that $\sum^{\infty}_{k=1}|\rho_{k}|<\infty$, $\sum^{\infty}_{k=1}|\sigma_{k}|<\infty$ and
$g(z)=\sum_{k}c_{k}\prod^{k-1}_{j=0}H(z+j).$ (56.01)
Then a solution of
$\sum_{k}g_{k}(-i)^{k}y^{(k)}(x)=V(x)y(x)\textrm{, }V(x)=\sum^{\infty}_{n=0}c_{n}e^{-inx},$
is such that
$\widehat{y}(z)=C^{z}\Gamma(z)^{m}e^{-R_{1}(z)}\prod^{\infty}_{j=1}\frac{\Gamma(z-\rho_{j})}{\Gamma(z-\sigma_{j})}.$
Also it holds
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=H(z)\Leftrightarrow\widehat{y}(z)=\prod^{\infty}_{n=0}\frac{1}{H(z+n)}$ (56.02)
and
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}C^{z}\Gamma(z)^{m}e^{-R_{1}(z)}\left(\prod^{\infty}_{j=1}\frac{\Gamma(z-\rho_{j})}{\Gamma(z-\sigma_{j})}\right)e^{izx}dz.$
III. Note that if $g(z)$ is 11-periodic, then (see Theorem 10 below) $H(z)$ is also 11-periodic. Hence
$g(z)=\sum_{k}c_{k}(H(z))^{k}.$ (56.03)
Set
$V_{1}(z)=\sum_{k}c_{k}z^{k},$ (56.04)
then
$H(z)=V^{(-1)}_{1}\left(g(z)\right).$ (56.05)
Hence given the $c_{k}$ of the potential $V(z)$ and $g_{n}$ of $g(z)$ we get, if $g(z)$ is 11-periodic
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=V_{1}^{(-1)}(g(z)).$ (56.06)
For example if $g(z)=\cos(2\pi z)$ and $V(z)=\sum^{\infty}_{n=1}\frac{e^{-inz}}{n}$, then
$\cos(2\pi z)=\sum^{\infty}_{n=1}\frac{H(z)^{n}}{n}=-\log(1-H(z)).$
Hence $H(z)=1-e^{-\cos(2\pi z)}$ and
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=1-e^{-\cos(2\pi z)}.$ (56.1)
However we can not solve with iteration this equation because it is 11-periodic and
$\widehat{y}(z)=\prod^{\infty}_{n=0}\frac{1}{1-e^{-\cos(2\pi(z+n))}}=\prod^{\infty}_{n=0}\frac{1}{1-e^{-\cos(2\pi z)}}.$
Moreover if $H(z)$ is semi-periodic i.e. exist $\lambda$ such that $\lambda\neq\pm 1$ and $0<|\lambda|\leq 1$, with
$H(z+1)=\lambda H(z)\textrm{, }\forall z\in D.$
Then
$g(z)=\sum_{k}c_{k}H(z)^{k}\lambda^{k(k-1)/2}.$
Hence
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=H(z).$
Moreover than this, if we assume that $\lambda(z)$ is semi-periodic and
$\lambda(z+1)=\xi\lambda(z)\textrm{, }0<|\xi|\leq 1\textrm{, }\xi\neq\pm 1.$
Then if
$H(z+1)=\lambda(z)H(z),$
we get easily
$H(z+n)=\xi^{n(n-1)/2}\lambda(z)^{n}H(z).$
Also if
$\widehat{y}(z+1)=H(z)\widehat{y}(z),$ (e1)
then
$\widehat{y}(z+n)=\xi^{n(n-1)(n-2)/6}\left(\lambda(z)\right)^{n(n-1)/2}H(z)^{n}\widehat{y}(z).$ (e2)
But as someone can see if we take $\widehat{y}(z)$ as
$\widehat{y}(z)=\sum^{\infty}_{k=-\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}H(z)^{-k},$ (e3)
then $(e1)$ holds if
$A_{k-1}(z+1)=A_{k}(z).$ (e4)
Hence $(e2)$ will holds also, since from $(e3)$ we can easily see that
$\widehat{y}(z+1)=H(z)\sum^{\infty}_{k=-\infty}A_{k-1}(z+1)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}H(z)^{-k}\Leftrightarrow$
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=H(z).$
Also we can write from $(e2)$:
$\sum^{\infty}_{k=-\infty}c_{k}\widehat{y}(z+k)=g(z)\widehat{y}(z)$
iff
$g(z)=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(z)^{k(k-1)/2}H(z)^{k}.$ (e5)
Theorem 9.0
Assume the function $\widehat{y}(z)$. If exists $A_{k}(z)$ such that
$\widehat{y}(z)=\sum_{k=-\infty}^{\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}H(z)^{-k}$
and
$A_{k}(z)=A_{k-1}(z+1),$
for these $\lambda(z),H(z)$, we will have
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=H(z)$
and the equation
$\sum^{\infty}_{k=-\infty}c_{k}\widehat{y}(z+k)=g(z)\widehat{y}(z)$
will be equivalent to
$g(z)=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(z)^{k(k-1)/2}H(z)^{k}.$
Remarks.
I.
The functions $\lambda(x)$ and $H(x)$, are such that
$\lambda(x+1)=\xi\lambda(x)\textrm{, }0<|\xi|\leq 1\textrm{, }\xi\neq\pm 1,$
$H(x+1)=\lambda(x)H(x)$
and
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}.$
(For example a possible construction of $\lambda(x)$ function, is to define a function $L(y,x)$ such that at a specific point $y=\xi$ (or maybe all $y\in D$), we have $L(\xi,x+1)=L(\xi,x)$, $\forall x$ and $\lambda(x)=\xi^{x}L(\xi,x)$, then $\lambda(x+1)=\xi^{x+1}L(\xi,x+1)=$ $\xi\left(\xi^{x}L(\xi,x)\right)$ $\xi\lambda(x)$.)
II. Hence if the function $g(x)=\sum^{\infty}_{n=0}g_{n}x^{n}$, which is related with the problem
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}y^{(n)}(x)=V(x)y(x),$
(where $V(x)$ is a continuous $2\pi$-periodic potential), is of the form
$g(x)=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(x)^{k(k-1)/2}H(x)^{k},$
then we have solution (of the problem) $y(x)$ such that
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=H(z)$
and
$\widehat{y}(z)=\sum_{k=-\infty}^{\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}H(z)^{-k},$
where $A_{k}(z)=A_{k-1}(z+1)$.
III. Hence for every given function $g(x)$ and sequence $c_{k}$ we have an expansion of the form (here for simplicity we take $V(x)$ to be a trigonometric polynomial of degree $N$):
$g(x)=\sum^{N}_{k=-N}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(x)^{k(k-1)/2}H(x)^{k},$
where $\lambda(x)$, $H(x)$ are defined as in I of these remarks. Hence
$g(x)=\sum^{N}_{k=-N}c_{k}\xi^{k(k-1)(k-2)/6}\left(\frac{H(x+1)}{H(x)}\right)^{k(k-1)/2}H(x)^{k}.$
Hence we have the polynomial equation
$S_{N}\left(\xi,\frac{H(x+1)}{H(x)},H(x)\right)=g(x),$ (e6)
where
$S_{N}(\xi,x,y)=\sum^{N}_{k=-N}c_{k}\xi^{k(k-1)(k-2)/6}x^{k(k-1)/2}y^{k}.$ (e7)
Hence solving (e6), with respect to $\lambda(x)=H(x+1)/H(x)$, we have
$H(x+1)=H(x)R_{N}(\xi,g(x),H(x)).$ (e8)
Hence
$\lambda(x)=R_{N}(\xi,g(x),H(x)),$ (e9)
where $R_{N}(\xi,x,y)$ is an algebraic function solution of (e7), with respect to $x$.
IV. Hence assuming that $g(x)$ is of the form:
$g(x)=\sum^{\infty}_{k=0}b_{k}\lambda(x)^{k},$
we have
$g(x)=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(x)^{k(k-1)/2}H(x)^{k},$
$\sum^{\infty}_{k=0}b_{k}x^{k}=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}x^{k(k-1)/2}H\left(\lambda^{(-1)}(x)\right)^{k}.$
Hence exists function $f(x)$ (by inversion) such that ($f(x)$ is known function)
$H\left(\lambda^{(-1)}(x)\right)=f(x).$
Hence
$x=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(x)^{k(k-1)/2}f(\lambda(x))^{k}\right).$
Hence
$\lambda^{(-1)}(x)=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}x^{k(k-1)/2}f(x)^{k}\right).$
Hence $\lambda(x)$ follows. Also then
$H(x)=f\left(\lambda(x)\right),$
and $H(x)$ also follows.
Note that
$\lambda(z+1)=\xi\lambda(z)\Leftrightarrow\lambda^{(-1)}(\xi z)-\lambda^{(-1)}(z)=1\Leftrightarrow$
$\lambda^{(-1)}\left(\xi^{z+1}\right)-\lambda^{(-1)}\left(\xi^{z}\right)=1.$ (a)
From relation $(a)$ we get that: for the function
$f(z)=H\left(\lambda^{(-1)}(z)\right),$ (b)
holds
$f\left(\xi^{z+1}\right)=H\left(\lambda^{(-1)}\left(\xi^{z+1}\right)\right)=H\left(1+\lambda^{(-1)}\left(\xi^{z}\right)\right)=$
$\lambda\left(\lambda^{(-1)}\left(\xi^{z}\right)\right)H\left(\lambda^{(-1)}\left(\xi^{z}\right)\right)=\xi^{z}f\left(\xi^{z}\right).$
Hence
$f(\xi z)=zf(z)$ (c)
and
$\lambda^{(-1)}(z)=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right),$ (d)
when the $g(z)$ and $c_{k}$ are given.
It holds from $(a),(d)$:
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}(\xi z)^{k(k-1)/2}f(\xi z)^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)=1$ (e)
Hence we have from $(c)$ and $(d)$:
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}(\xi z)^{k(k-1)/2}\left(zf(z)\right)^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)=1.$ (d1)
Note that we have an identity also:
$g^{(-1)}\left(f(z)^{-1}\sum^{\infty}_{k=-\infty}c_{k-1}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)=$
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi z\right)^{k(k-1)/2}\left(zf(z)\right)^{k}\right).$ (d2)
Hence
$g^{(-1)}\left(f(z)^{-1}\sum^{\infty}_{k=-\infty}c_{k-1}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)=1.$ (d3)
Hence from $(d2),(d3)$, we get
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi z\right)^{k(k-1)/2}\left(zf(z)\right)^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)=1.$ (d4)
Hence equation $(d3)$ is of the form
$g^{(-1)}\left(f(z)^{-1}F\left(z,f(z)\right)\right)-g^{(-1)}\left(G\left(z,f(z)\right)\right)=1,$ (d5)
where $F,G,g$ are the known functions:
$F(z,w)=\sum^{\infty}_{k=-\infty}c_{k-1}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}w^{k}$
and
$G(z,w)=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}w^{k},$
(since they have known coeficients $c_{k}$).
Hence assuming $\lambda(z+1)=\xi\lambda(z)$, we get from $H(z+1)=\lambda(z)H(z)$ that if $f(z)$ is $f(z)=H\left(\lambda^{(-1)}(z)\right)$, then $f(\xi z)=zf(z)$. Also $(d5)$ is an ordinary (non-differential) transcendental equation with unkown function the $f(z)$. Hence we have the next
Theorem 9.1
If for the functions $F(z,w),G(z,w)$ defined above, we have
$g^{(-1)}\left(H(z)^{-1}F\left(\lambda(z),H(z)\right)\right)-g^{(-1)}\left(G\left(\lambda(z),H(z)\right)\right)=1,$
where $\lambda(z+1)=\xi\lambda(z)$ and $H(z+1)=\lambda(z)H(z)$, then it holds
$\sum^{\infty}_{k=-\infty}c_{k}\widehat{y}(z+k)=g(z)\widehat{y}(z),$
when
$\frac{\widehat{y}(z+1)}{\widehat{y}(z)}=H(z)$
and
$g(z)=\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\lambda(z)^{k(k-1)/2}H(z)^{k}.$
Then we have
$\widehat{y}(z)=\sum^{\infty}_{k=-\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}H(z)^{-k},$
where
$A_{n-1}(z+1)=A_{n}(z).$
Remark.
If the function $e\left(g^{(-1)}(z)\right)$ is one to one, then
$H(z)^{-1}F\left(\lambda(z),H(z)\right)=G\left(\lambda(z),H(z)\right)$
or equivalently
$f(z)^{-1}F\left(z,f(z)\right)=G\left(z,f(z)\right).$
Theorem 9.2
The differential equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x),$
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx},$
have solution
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\widehat{y}(z)e^{izx}dz,$
where
$\widehat{y}(z)=\sum^{\infty}_{k=-\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}f\left(\lambda(z)\right)^{-k},$
with $A_{n-1}(z+1)=A_{n}(z)$ and $\lambda(z+1)=\xi\lambda(z)$. The equations of finding $f(z)$ are both
$g^{(-1)}\left(f(z)^{-1}\sum^{\infty}_{k=-\infty}c_{k-1}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right)=1$ (d6)
and
$f(\xi z)=zf(z).$ (d7)
Note that
$\lambda^{(-1)}(z)=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right).$
Remarks.
Set now
$C(z):=\int^{\infty}_{-\infty}H(t)\lambda(t)^{z}dt\textrm{, }z\in\textbf{C}.$
Then
$C(z+1)=\int^{\infty}_{-\infty}H(t)\lambda(t)^{z+1}dt=\int^{\infty}_{-\infty}H(t)\lambda(t)\lambda(t)^{z}dt=$
$\int^{\infty}_{-\infty}H(t+1)\lambda(t)^{z}dt=\int^{\infty}_{-\infty}H(t)\lambda(t-1)^{z}dt=$
$\int^{\infty}_{-\infty}H(t)\left(\xi^{-1}\lambda(t)\right)^{z}dt=\xi^{-z}\int^{\infty}_{-\infty}H(t)\lambda(t)^{z}dt=\xi^{-z}C(z).$
Hence
$C(z+1)=\xi^{-z}C(z)$
and
$C(z)=C_{0}(\xi_{1},z)\xi^{-z(z-1)/2}\textrm{, }z\in\textbf{C},$
$\lambda(z)=\xi^{z}\lambda_{0}(\xi_{2},z)\textrm{, }\lambda(z+1)=\xi\lambda(z),$
where $C_{0}(\xi_{1},z)$ and $\lambda_{0}(\xi_{2},z)$ are 11-periodic i.e. $C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z)$ and $\lambda_{0}(\xi_{2},z+1)=\lambda_{0}(\xi_{2},z)$. Hence if we set
$\Theta(z):=\sum_{n\in\textbf{\scriptsize Z\normalsize}}C(n)\lambda(z)^{-n},$
then
$\Theta(z+1)=\sum_{n\in\textbf{\scriptsize Z\normalsize}}C(n)\lambda(z+1)^{-n}=\sum_{n\in\textbf{\scriptsize Z\normalsize}}C(n)\left(\xi\lambda(z)\right)^{-n}=$
$\lambda(z)\sum_{n\in\textbf{\scriptsize Z\normalsize}}C(n+1)\lambda(z)^{-n-1}=\lambda(z)\Theta(z).$
Hence there exists constant $c_{0}$ such that
$\Theta(z)=c_{0}\sum^{\infty}_{n=-\infty}\xi^{-n(n-1)/2}\lambda(z)^{-n}=c_{0}\sum^{\infty}_{n=-\infty}\xi^{-n(n+1)/2}\lambda(z)^{n}$
and
$\Theta(z+1)=\lambda(z)\Theta(z).$
Set
$G(z):=\int^{\infty}_{-\infty}C(t)\left(\lambda(z)\right)^{-t}dt\textrm{, }z\in\textbf{C}.$
Then
$G(z+1)=\int^{\infty}_{-\infty}C(t)\left(\lambda(z+1)\right)^{-t}dt=\int^{\infty}_{-\infty}C(t)\left(\xi\lambda(z)\right)^{-t}dt=$
$\int^{\infty}_{-\infty}C(t)\xi^{-t}\left(\lambda(z)\right)^{-t}dt=\int^{\infty}_{-\infty}C(t+1)\left(\lambda(z)\right)^{-t}dt=$
$\lambda(z)\int^{\infty}_{-\infty}C(t+1)\left(\lambda(z)\right)^{-t-1}dt=\lambda(z)G(z).$
Hence again
$G(z+1)=\lambda(z)G(z).$
Now as one can see $(d7)$ can be writen in the form $f\left(\xi^{z+1}\right)=\xi^{z}f\left(\xi^{z}\right)$. Hence exists $C(z)$ such that $f\left(\xi^{z}\right)=C(z)^{-1}$. Hence $(d6)$ becomes
$g^{(-1)}\left(f(\xi^{z})^{-1}\sum^{\infty}_{k=-\infty}c_{k-1}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}f(\xi^{z})^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}f(\xi^{z})^{k}\right)=1.$
Hence
$g^{(-1)}\left(C(z)\sum^{\infty}_{k=-\infty}c_{k-1}\xi^{k(k-1)(k-2)/6+zk(k-1)/2}C(z)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6+zk(k-1)/2}C(z)^{-k}\right)=1.$
Hence
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k+1)/6+zk(k+1)/2}C(z)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6+zk(k-1)/2}C(z)^{-k}\right)=1\Leftrightarrow$
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{k(k-1)/2}\xi^{zk(k+1)/2}C(z)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}C(z)^{-k}\right)=1\Leftrightarrow$
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{k(k-1)/2}\xi^{zk(k+1)/2}\xi^{-kz}C(z+1)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}C(z)^{-k}\right)=1\Leftrightarrow$
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{(z+1)k(k-1)/2}C(z+1)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}C(z)^{-k}\right)=1.$
Also if in relation $(d1)$ we set $z\rightarrow\xi^{z}$, then
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi^{z+1}\right)^{k(k-1)/2}\left(\xi^{z}f(\xi^{z})\right)^{k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi^{z}\right)^{k(k-1)/2}f(\xi^{z})^{k}\right)=1\Leftrightarrow$
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi^{z+1}\right)^{k(k-1)/2}C(z+1)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi^{z}\right)^{k(k-1)/2}C(z)^{-k}\right)=1.$
If we set
$A(z):=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\left(\xi^{z}\right)^{k(k-1)/2}C(z)^{-k}\right),$
then
A(z+1)A(z)=1A(z+1)-A(z)=1
and the function $\xi^{A(z)}$ is semiperiodic function, with period 1 and weight $\xi$. Hence
$\xi^{A(z)}=\xi^{z}\mu_{0}(z),$
where $\mu_{0}(z)$ is 11-periodic function. Hence equivalent
$A(z)=z+\log_{\xi}\left(\mu_{0}(z)\right).$
Hence we get the next
Theorem 9.3
The two equations $(d6),(d7)$ of Theorem 9.2, can be writen as:
$\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f\left(z\right)^{k}=g\left(\log_{\xi}(z)+\log_{\xi}\left(\mu_{0}\left(\log_{\xi}(z)\right)\right)\right),$
where $\mu_{0}(z)$ is 11-periodic function ($\mu_{0}(z+1)=\mu_{0}(z)$).
And
$f(z)=C\left(\log_{\xi}z\right)^{-1},$
with $C(z+1)=\xi^{-z}C(z)$, (hence $C(z)=\xi^{-z(z-1)/2}C_{0}(\xi_{1},z)$, and $C_{0}(\xi_{1},z)$ is 11-periodic).
From Theorem 9.3 and Theorem 9.1, we have
Theorem 9.4
Assume the problem:
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x)\textrm{, }g(x)=\sum^{\infty}_{n=0}g_{n}x^{n},$
with periodic potential
$V(x)=\sum^{\infty}_{n=-\infty}c_{n}e^{-inx}.$
Then if
$f(z)=C\left(\log_{\xi}z\right)^{-1},$
with $C(z)=\xi^{-z(z-1)/2}C_{0}(\xi_{1},z)$ and $C_{0}(\xi_{1},z)$ is 11-periodic function such that
$\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}\xi^{kz(z-1)/2}C_{0}\left(\xi_{1},z\right)^{-k}=g\left(z+\log_{\xi}\left(\mu_{0}\left(z\right)\right)\right),$
where $\mu_{0}(z)$ is also any 11-periodic function. Then function (solution) $y(x)$ (of the above problem) is such that
$\widehat{y}(z)=\sum^{\infty}_{k=-\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}f\left(\lambda(z)\right)^{-k},$
where $A_{n-1}(z+1)=A_{n}(z)$ and
$\lambda^{(-1)}(z)=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}z^{k(k-1)/2}f(z)^{k}\right).$
Therorem 9.5
The two equations $(d6),(d7)$ of Theorem 9.1, can be writen as:
If
$f(z)=C\left(\log_{\xi}z\right)^{-1}$
and
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{(z+1)k(k-1)/2}C(z+1)^{-k}\right)$-
$g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}C(z)^{-k}\right)=1,$
where $C(z)$ is such that $C(z+1)=\xi^{-z}C(z)$, then
$\widehat{y}(z)=\sum^{\infty}_{k=-\infty}A_{k}(z)\xi^{-k(k-1)(k-2)/6}\lambda(z)^{-k(k-1)/2}f\left(\lambda(z)\right)^{-k},$
with $A_{n-1}(z+1)=A_{n}(z)$ and $\lambda(z+1)=\xi\lambda(z)$.
Set now
$F(z,w):=g^{(-1)}\left(\sum^{\infty}_{k=-\infty}c_{k}\xi^{k(k-1)(k-2)/6}\xi^{zk(k-1)/2}\xi^{kz(z-1)/2}w^{-k}\right),$ (d8)
where the $c_{k}$ are given from the potential $V(z)$ and $g(z)$ are given from the DE’s of Theorems 9.1-9.5. We can write also $C(z+1)=\xi^{-z}C(z)$, where $C(z)=\xi^{-z(z-1)/2}C_{0}(\xi_{1},z)$ and $C_{0}(\xi_{1},z)$ is an 11-periodic function. Then we have
$F\left(z+1,C_{0}(\xi_{1},z+1)\right)-F\left(z,C_{0}(\xi_{1},z)\right)=1$ (a)
and
$C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z).$ (b)
Set
$P(z,w):=\exp\left(2\pi iF(z,w)\right)$ (c)
and
$G(z):=\frac{1}{C(z)}.$ (c0)
Then
$G(z+1)=\frac{1}{C(z+1)}=\xi^{z}G(z),$ (c1)
where $G(z):=\xi^{z(z-1)/2}G_{0}(\xi_{1},z)$ and $G_{0}(\xi_{1},z)$ is any 11-periodic function. Then from $(a)$ we have
$P(z+1,C_{0}(\xi_{1},z+1))=\exp\left(2\pi iF(z+1,C_{0}(\xi_{1},z+1))\right)=$
$\exp\left(2\pi iF(z,C_{0}(\xi_{1},z))\right)=P(z,C_{0}(\xi_{1},z)).$ (c2)
Also from $(d8)$, when $\xi^{\tau}=e^{2\pi i\tau}$, we have
$F(z+1,w)=F(z,\xi^{-z}w)\textrm{, }(z,w)\in D\times S.$ (c3)
Hence
$F(z+1,G(z+1))=F(z,\xi^{-z}G(z+1))=F(z,G(z)).$ (c4)
By this way we have
$P(z+1,G(z+1))=P(z,G(z))$ (c5)
and
$P(z+1,C_{0}(\xi_{1},z+1))=P(z,C_{0}(\xi_{1},z)).$ (c6)
Also
$P(z+1,w)=P(z,\xi^{-z}w).$ (c7)
Hence we get the next:
Theorem 9.5.1
Assume that equation
$P(z+1,Y(z))=P(z,X(z))\textrm{, }z\in D,$ (c8)
have equivalent solution
$Y(z)=S(z,X(z))\textrm{, }z\in D.$ (c9)
Then we can write that: Equation
$P(z+1,C_{0}(\xi_{1},z+1))=P(z,C_{0}(\xi_{1},z))\textrm{, }z\in D,$ (g1)
have equivalent unique solution
$C_{0}(\xi_{1},z+1)=S(z,C_{0}(\xi_{1},z))\textrm{, }z\in D.$ (g2)
Then all following three results are equivalent to each other:
i)
$P(z+1,C_{0}(\xi_{1},z+1))=P(z,C_{0}(\xi_{1},z))$
and $C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z)$.
ii)
$C_{0}(\xi_{1},z+1)=S(z,C_{0}(\xi_{1},z))$
and $C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z)$.
iii)
$C_{0}(\xi_{1},z)=S(z,C_{0}(\xi_{1},z))$
and
$S(z+1,C_{0}(\xi_{1},z+1))=S(z,C_{0}(\xi_{1},z+1)).$
iv)
$C_{0}(\xi_{1},z)=S(z,C_{0}(\xi_{1},z))$
and
$S(z+1,C_{0}(\xi_{1},z))=S(z,C_{0}(\xi_{1},z)).$
Proof.
Case $\textbf{(i)}\Leftrightarrow\textbf{(ii)}$ is obvious.
The case $\textbf{(ii)}\Rightarrow\textbf{(iii)}$ is
$C_{0}(\xi_{1},z)=C_{0}(\xi_{1},z+1)=S(z,C_{0}(\xi_{1},z))\Rightarrow$
$C_{0}(\xi_{1},z+1)=S(z+1,C_{0}(\xi_{1},z+1))=S(z,C_{0}(\xi_{1},z+1)).$
The case $\textbf{(iii)}\Rightarrow\textbf{(ii)}$ is
$C_{0}(\xi_{1},z)=S(z,C_{0}(\xi_{1},z))\Rightarrow C_{0}(\xi_{1},z+1)=S(z+1,C_{0}(\xi_{1},z+1))=$
$S(z,C_{0}(\xi_{1},z+1)).$
Hence the equation $X(z)=S(z,X(z))$ have two solutions $C_{0}(\xi_{1},z)$ and $C_{0}(\xi_{1},z+1)$. Hence it must be $C_{0}(\xi_{1},z)=C_{0}(\xi_{1},z+1)$.
The case $\textbf{(iv)}\Rightarrow\textbf{(iii)}$ is:
$S(z+1,C_{0}(\xi_{1},z))=S(z,C_{0}(\xi_{1},z))=C_{0}(\xi_{1},z)$
Hence
$C_{0}(\xi_{1},z+1)=S(z+1,C_{0}(\xi_{1},z+1))$
and
$S(z+1,C_{0}(\xi_{1},z))=C_{0}(\xi_{1},z).$
Hence $C_{0}(\xi_{1},z)=C_{0}(\xi_{1},z+1)$.
The case $\textbf{(iii)}\Rightarrow\textbf{(iv)}$ is
$C_{0}(\xi_{1},z+1)=S(z+1,C_{0}(\xi_{1},z+1))=S(z,C_{0}(\xi_{1},z+1))$
and
$C_{0}(\xi_{1},z)=S(z,C_{0}(\xi_{1},z)).$
Hence
$C_{0}(\xi_{1},z)=C_{0}(\xi_{1},z+1).$
QED
Theorem 9.5.2
The equations $(a),(d8),(b)$, that solve the problem of Theorems 9.1-9.5 are equivalent to $(c2),(c),(b)$ and hence to (from Theorem 9.5.1):
i)
$C_{0}(\xi_{1},z)=S(z,C_{0}(\xi_{1},z)),$ (g3)
$C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z),$ (g4)
where $S(z,w)$ is defined from $(c8),(c9)$.
Moreover then
ii)
$S(z,w)\textrm{ is equivalent to }P(z,w)=\exp(2\pi iF(z,w)),$ (g5)
where the function $F(z,w)$ is given from $(d8)$. The equivalence means that function $C_{0}(\xi_{1},z)$ is solution of ”almost” the same equations i.e.
$P(z+1,C_{0}(\xi_{1},z+1))=P(z,C_{0}(\xi_{1},z))$
and
$S(z+1,C_{0}(\xi_{1},z+1))=S(z,C_{0}(\xi_{1},z)).$
iii)
$C_{0}(\xi_{1},z)=P(z,C_{0}(\xi_{1},z)),$ (g6)
$C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z).$ (g7)
Proof.
Note that from Theorem 9.5.1 we have
$P(z+1,C_{0}(\xi_{1},z+1))=P(z,C_{0}(\xi_{1},z))$
and $C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z)$. Also then equivalent
$S(z+1,C_{0}(\xi_{1},z+1))=S(z,C_{0}(\xi_{1},z))$
and $C_{0}(\xi_{1},z+1)=C_{0}(\xi_{1},z)$. Hence we can say that $P(z,w)$ and $S(z,w)$ have the same ”shape”.
IV. If we assume the equation $y^{\prime\prime}(x)=V(x)y(x)$, where $V(x)=\sum^{N}_{n=0}c_{n}e^{-inx}$, is a trigonometric polynomial, then under certain conditions we have
$\sum^{N}_{n=0}c_{n}\widehat{y}(z+n)=-z^{2}\widehat{y}(z).$
This equation is a simple special functions functional equation and we can assume is taking known solutions. For example, with $N=2$ the equation
$c_{0}f(x)+c_{1}f(x+1)+c_{2}f(x+2)=-x^{2}f(x)$
gives a solution $f(x)$ and hence the equation
$y^{\prime\prime}(x)=\left(\sum^{2}_{n=0}c_{n}e^{-inx}\right)y(x),$
have solution $y(x)$ such that $\widehat{y}(x)=f(x)\Leftrightarrow y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}f(t)e^{itx}dt$.
More general if $g_{n}=g^{(n)}(0)/n!$, the equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(x)=V(x)y(x),$ (56.0a)
where the potential $V(x)$ is of the form
$V(x)=\sum_{n=0}^{N}c_{n}e^{-inx},$
have solution $y(x)$ such that
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}f(t)e^{itx}dt,$
with $f(x)$ solution of the following functional equation
$c_{0}f(x)+c_{1}f(x+1)+c_{2}f(x+2)+\ldots+c_{N}f(x+N)=g(x)f(x)$
and
$g(x)=\sum^{\infty}_{n=0}g_{n}x^{n}.$
Hence if $y^{\prime\prime}(x)=V(x)y(x)$, $V(x)=\sum^{N}_{n=0}c_{n}e^{-inx}$ and $f(x)$ is such that
$c_{0}f(x)+c_{1}f(x+1)+\ldots+c_{N}f(x+N)=-x^{2}f(x),$
then
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}f(t)e^{itx}dt.$
V. If $g(z)$ is of the form
$g(z)=1+\frac{a_{0}(z)}{1+}\frac{-a_{1}(z)}{(1+a_{1}(z))+}\frac{-a_{2}(z)}{(1+a_{2}(z))+}\frac{-a_{3}(z)}{(1+a_{3}(z))+}\ldots=$
$1+\sum^{\infty}_{k=1}\prod^{k-1}_{j=0}a_{j}(z),$
where $a_{j}(z)=\frac{A}{B}c_{j}^{*}H(z+j)$, then
$g(z)=1+\frac{c_{0}^{*}H(z)}{1+}\frac{-\frac{A}{B}c_{1}^{*}H(z+1)}{\left(1+\frac{A}{B}c_{1}^{*}H(z+1)\right)+}\frac{-\left(\frac{A}{B}\right)^{2}c_{2}^{*}H(z+2)}{\left(1+\left(\frac{A}{B}\right)^{2}c_{2}^{*}H(z+2)\right)+}\ldots.$ (cf)
Hence given $g(z)$, we expand it in the form $(cf)$ and find $H(z)$. Then using Weierstrass theorem
$H(z)=Cz^{m}e^{-R(z)}\frac{P(z)}{Q(z)},$
or (56.00), we find the solution
$\widehat{y}(z)=\left(\frac{A}{B}\right)^{z}C^{z}\Gamma(z)^{m}e^{-R_{1}(z)}\prod^{\infty}_{j=1}\frac{\Gamma(z-\rho_{j})}{\Gamma(z-\sigma_{j})}.$
Theorem 9.6
Assume that
$f(z,w)=\sum^{\infty}_{n=-\infty}A_{n}(z)w^{-n}.$
i) If $A_{n}(z)$ satisfies the relation
$A_{n-1}(z+1)=A_{n}(z),$ (f)
then equivalent we have that
$f\left(z+1,w\right)=wf\left(z,w\right).$
ii) Also if
$\widehat{y}(z)=\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}$
and holds $(f)$ and $\lambda(z+1)=\xi\lambda(z)$, then:
$H(z+1)=\lambda(z)H(z)$
is equivalent to
$\widehat{y}(z+1)=H(z)\widehat{y}(z).$
Proof.
The relation $A_{n-1}(z+1)=A_{n}(z)$ is equivalent to
$\sum^{\infty}_{n=-\infty}A_{n-1}(z+1)w^{-n}=\sum^{\infty}_{n=-\infty}A_{n}(z)w^{-n}\Leftrightarrow$
$\sum^{\infty}_{n=-\infty}A_{n-1}(z+1)w^{-n}=w\sum^{\infty}_{n=-\infty}A_{n}(z)w^{-n}\Leftrightarrow$
$f\left(z+1,w\right)=wf(z,w).$
For the other statements, we have
$\widehat{y}(z+1)=$
$\sum^{\infty}_{n=-\infty}A_{n}(z+1)\xi^{-n(n-1)(n-2)/6}\xi^{-n(n-1)/2}\lambda(z)^{-n(n-1)/2}\lambda(z)^{-n}H(z+1)^{-n}=$
$\sum^{\infty}_{n=-\infty}A_{n}(z+1)\xi^{-n(n-1)(n-2)/6}\xi^{-n(n-1)/2}\lambda(z)^{-n(n-1)/2}\lambda(z)^{-n}H(z)^{-n}=$
$\sum^{\infty}_{n=-\infty}A_{n}(z+1)\xi^{-n(n-1)(n+1)/6}\lambda(z)^{-n(n+1)/2}H(z)^{-n}=$
$H(z)\sum^{\infty}_{n=-\infty}A_{n-1}(z+1)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}=$
$H(z)\widehat{y}(z).$
Assuming that $\widehat{y}(z+1)=H(z)\widehat{y}(z)$, we have
$\sum_{n}A_{n}(z+1)\xi^{-n(n-1)(n-2)/6}\lambda(z+1)^{-n(n-1)/2}H(z+1)^{-n}=$
$H(z)\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}\Leftrightarrow$
$\sum_{n}A_{n}(z+1)\xi^{-n(n-1)(n-2)/6}\xi^{-n(n-1)/2}\lambda(z)^{-n(n-1)/2}H(z+1)^{-n}=$
$H(z)\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}\Leftrightarrow$
$\sum_{n}A_{n}(z+1)\xi^{-n(n-1)(n+1)/2}\lambda(z)^{-n(n+1)/2}\lambda(z)^{n}H(z+1)^{-n}=$
$H(z)\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}\Leftrightarrow$
$\sum_{n}A_{n-1}(z+1)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}\lambda(z)^{n-1}H(z+1)^{-n+1}=$
$H(z)\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}\Leftrightarrow$
$\left(\lambda(z)^{-1}H(z+1)\right)\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}\left(\lambda(z)^{-1}H(z+1)\right)^{-n}=$
$H(z)\sum_{n}A_{n}(z)\xi^{-n(n-1)(n-2)/6}\lambda(z)^{-n(n-1)/2}H(z)^{-n}.$
Hence obviously
$H(z+1)=\lambda(z)H(z).$
Theorem 9.7
With the notation of Theorem 9.0a we have
$H(z+j)=\left(\prod^{j-1}_{k=0}\lambda(z+k)\right)H(z)$
and
$\widehat{y}(z+n)=\left(\prod^{n-1}_{j=0}\prod^{j-1}_{k=0}\lambda(z+k)\right)H(z)^{n}\widehat{y}(z).$
Hence
$\sum^{\infty}_{n=0}c_{n}\widehat{y}(z+n)=g(z)\widehat{y}(z)$
and
$g(z)=\sum^{\infty}_{n=0}c_{n}\left(\prod^{n-1}_{j=0}\prod^{j-1}_{k=0}\lambda(z+k)\right)H(z)^{n}.$
Also then
$\widehat{y}(z)=\sum^{\infty}_{n=-\infty}A_{n}(z)H(z)^{n},$
where
$A_{n+1}(z+1)\lambda(z)^{n+1}=A_{n}(z).$
Theorem 9.8
$g(z)=\sum_{n}c_{n}\prod^{n-1}_{j=0}H\left(z+\frac{2\pi j}{T}\right),$
then if $H(z)=\exp\left(H_{1}(z)\right)$, we have
$g(z)=\sum_{n}c_{n}\exp\left(\sum^{n-1}_{j=0}H_{1}\left(z+\frac{2\pi j}{T}\right)\right).$
Hence assuming the equation
$\sum^{n-1}_{j=0}H_{1,n}\left(z+\frac{2\pi j}{T}\right)=g(z)H_{1,n}(z),$
we have a solution
$\sum^{n-1}_{j=0}f_{n}\left(z+\frac{2\pi j}{T}\right)=g(z)f_{n}(z)$
and $f_{n}(z)=H_{1,n}(z)$ and if $\lim_{n\rightarrow\infty}H_{1,n}(z)=\lim_{n\rightarrow\infty}f_{n}(z)=f(z)$. Hence
$\sum^{\infty}_{j=0}f\left(z+\frac{2\pi j}{T}\right)=g(z)f(z).$
Theorem 9.9
Assume
$V(z)=\sum^{\infty}_{n=0}c_{n}e^{-inz}$
and
$H(z)=\frac{(z+A_{1})(z+A_{2})\ldots(z+A_{N})}{(z+B_{1})(z+B_{2})\ldots(z+B_{M})}e^{-az}.$
Then
$g(z)=\sum^{\infty}_{k=0}\left(\frac{A}{B}\right)^{k}c_{k}\frac{(z+A_{1})_{k}(z+A_{2})_{k}\ldots(z+A_{N})_{k}}{(z+B_{1})_{k}(z+B_{2})_{k}\ldots(z+B_{M})_{k}}\frac{e^{-akz}e^{-ak(k+1)/2}}{k!}.$
Proof.
Easy follows from (53) of Theorem 9.
If $f:\textbf{N}\rightarrow\textbf{N}$ is increasing function and $\chi:\textbf{N}\rightarrow\textbf{C}$, then the function
$\theta(q)=\sum_{n=1}^{\infty}\chi(n)q^{f(n)},$
can be written as Lambert series:
$\theta(q)=\sum^{\infty}_{n=1}\frac{B(n)q^{n}}{1-q^{n}},$
where
$B(n)=\sum_{f(d)|n}\chi(d)\mu\left(\frac{n}{f(d)}\right).$
Also then
$\theta(q)=\sum^{\infty}_{n=1}q^{n}\sum_{d|n}B(d)=\sum^{\infty}_{n=1}q^{n}A(n),$
where
$A(n)=\sum_{d|n}\sum_{f(\delta)|d}\chi(\delta)\mu\left(\frac{d}{f(\delta)}\right).$
If $A(n)$ is $T_{0}$-periodic, $T_{0}\in\textbf{N}$, $T_{0}>1$, then
$\theta(q)=\left(1-q^{T_{0}}\right)^{-1}\sum^{T_{0}}_{n=1}A(n)q^{n},$
is rational function of $q$.
Theorem 10.
Given $g(x)=\sum^{\infty}_{n=0}g_{n}x^{n}$ and $V(x)=\sum^{\infty}_{n=0}c_{n}e^{-2\pi inx/T}$, the equation
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}y^{(n)}(x)=V(x)y(x),$
is equivalent to
$\sum^{\infty}_{n=0}c_{n}\widehat{y}\left(z+\frac{2\pi}{T}n\right)=g(z)\widehat{y}(z).$
If $H(z)$ is defined from
$\widehat{y}\left(z+\frac{2\pi}{T}\right)=H(z)\widehat{y}(z),$
then it must hold
$g(z)=\sum_{n=0}^{\infty}c_{n}\prod^{n-1}_{j=0}H\left(z+\frac{2\pi}{T}j\right).$
Hence assuming that $H(z)$ is of the form
$H(z)=Cz^{m}e^{-R(z)}\theta(z),$
where $R(z)$ holomorphic function and $\theta(z)$ meromophic function such that
$\theta(z)=\sum_{n=1}^{\infty}\chi(n)z^{f(n)},$
with $f:\textbf{N}\rightarrow\textbf{N}$ is increasing function and $\chi:\textbf{N}\rightarrow\textbf{C}$ and the function
$A(n)=\sum_{d|n}\sum_{f(\delta)|d}\chi(\delta)\mu\left(\frac{d}{f(\delta)}\right),$
is $T_{0}$-periodic, $T_{0}\in\textbf{N}$, $T_{0}>1$. Then
$\theta(z)=\left(1-z^{T_{0}}\right)^{-1}\sum^{T_{0}}_{n=1}A(n)z^{n},$
is rational function of $z$. Hence
$y(x)=\frac{1}{2\pi}\int^{+\infty}_{-\infty}C^{z}\Gamma(z)^{m}e^{-R_{1}(z)}\left(\frac{\prod^{T_{0}}_{j=1}\Gamma(z-\rho_{j})}{\prod^{T_{0}}_{j=1}\Gamma(z-\sigma_{j})}\right)e^{izx}dz.$
Definition 10.1
We define the operator T such that if
$g(z)=\sum^{\infty}_{n=0}c_{n}\prod^{n-1}_{j=0}H(z+j),$
then
$\left(\textbf{T}g\right)(z)=\sum^{\infty}_{n=0}c_{n-1}\prod^{n-1}_{j=0}H(z+j),$
with $c_{-1}=0$ and $\prod^{-1}_{j=0}a_{j}=1$.
Theorem 10.2
$H(z)=\frac{\left(\textbf{T}g\right)(z)}{g(z+1)}.$
Proof.
We have easily
$g(z+1)=\sum^{\infty}_{n=0}c_{n}\prod^{n-1}_{j=0}H(z+j+1)=H(z)^{-1}\sum^{\infty}_{n=0}c_{n}\prod^{n}_{j=0}H(z+j)=$
$H(z)^{-1}\sum^{\infty}_{n=0}c_{n-1}\prod^{n-1}_{j=0}H(z+j)=H(z)^{-1}\left(\textbf{T}g\right)(z).$
Theorem 10.3
Assume that $e_{n}(z)$ is a base in A subset of R and $g^{*}_{n}=\left\langle g(z),e_{n}(z)\right\rangle$, where $g(z)=\sum^{\infty}_{n=0}g^{*}_{n}e_{n}(x)$. Also
$V(z)=\sum^{\infty}_{n=0}c_{n}e^{-inz}$
and
$\sum^{\infty}_{n=0}g_{n}(-i)^{n}\frac{d^{n}}{dx^{n}}y(z)=V(x)y(x).$
For to solve the above equation, set $J_{n,k}(a_{(.)})$ to be an operator such that:
$\exp\left(\sum^{\infty}_{k=0}a_{k}B_{k,n}(z)\right)=\sum^{\infty}_{k=0}J_{n,k}(a_{(.)})B_{k,n}(z),$
where
$B_{k,n}(z)=\sum^{k}_{l=0}C_{k,l}p_{k-l}(n)z^{l}\textrm{, }p_{\tau}(n)=\sum^{n-1}_{j=0}j^{\tau}.$
If $e_{n}(z)$ is a certain base, we write
$B_{k,n}(z)=\sum^{\infty}_{l=1}S_{k,n,l}e_{l}(z).$
and $S^{*}_{k,n,l}$ is the matrix inverse of $S_{k,n,l}$, in the sense
$\sum^{\infty}_{l=1}S_{k,n,l}S^{*}_{k^{\prime},n,l}=R_{n}\delta_{k^{\prime},k},$
where $\delta_{k^{\prime},k}=1$, if $k=k^{\prime}$ and $\delta_{k^{\prime},k}=0$, if $k^{\prime}\neq k$. Also we define $a^{\prime}_{n,k}$ to be the arithmetic inverse of $\sum_{d|n}c_{d}\mu(n/d)$, in the sense
$\sum_{d_{1}|A}\left(\sum_{d|d_{1}}c_{d}\mu(d_{1}/d)\right)a^{\prime}_{A/d_{1},B}=\delta_{A,1}\delta_{B,1}.$
Then with a suitable choise of $R_{n}$, we get
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\prod^{\infty}_{n=0}\exp\left(-\sum^{\infty}_{k=0}J_{n,k}^{(-1)}\left(G^{-1}\sum^{\infty}_{l=1}g^{*}_{l}\sum_{d_{1}|n,d_{2}|l}S^{*}_{k,d_{1},d_{2}}a^{\prime}_{n/d_{1},l/d_{2}}\right)(z+n)^{k}\right)e^{izx}dz.$
Proof.
We write
$g(z)=\sum^{\infty}_{n=0}c_{n}\prod^{n-1}_{j=0}H(z+j),$
where
$H(z)=\exp\left(\sum^{\infty}_{k=0}a_{k}z^{k}\right)$
Hence if $C_{n,k}$ denotes the binomial function i.e. $C_{n,k}=\frac{n!}{(n-k)!k!}$, then
$\prod^{n-1}_{j=0}H(z+j)=\exp\left(\sum^{\infty}_{k=0}a_{k}\sum^{n-1}_{j=0}(z+j)^{k}\right)=$
$\exp\left(\sum^{\infty}_{k=0}a_{k}\sum^{n-1}_{j=0}\sum^{k}_{l=0}C_{k,l}j^{k-l}z^{l}\right)=$
$\exp\left(\sum^{\infty}_{k=0}a_{k}\sum^{k}_{l=0}C_{k,l}\sum^{n-1}_{j=0}j^{k-l}z^{l}\right)=$
$\exp\left(\sum^{\infty}_{k=0}a_{k}\sum^{k}_{l=0}C_{k,l}p_{k-l}(n)z^{l}\right).$
Hence if we denote
$B_{k,n}(z):=\sum^{k}_{l=0}C_{k,l}p_{k-l}(n)z^{l},$
then
$g(z)=\sum^{\infty}_{n=0}c_{n}\exp\left(\sum^{\infty}_{k=0}a_{k}B_{k,n}(z)\right).$
Now we assume the transformation $J$ such that
$\exp\left(\sum^{\infty}_{k=0}a_{k}B_{k,n}(z)\right)=\sum^{\infty}_{k=0}J_{n,k}(a_{(.)})B_{k,n}(z).$
This equivalently give (if we denote $h_{n,k}=J_{n,k}(a_{(.)})$):
$\exp\left(\sum^{n-1}_{j=0}\sum^{\infty}_{k=0}a_{k}(z+j)^{k}\right)=\sum^{n-1}_{j=0}\sum^{\infty}_{k=0}h_{n,k}(z+j)^{k}\Leftrightarrow$
$\prod^{n-1}_{j=0}\exp\left(\sum^{\infty}_{k=0}a_{k}(z+j)^{k}\right)=\sum^{n-1}_{j=0}\log\left(\exp\left(\sum^{\infty}_{k=0}h_{n,k}(z+j)^{k}\right)\right)\Leftrightarrow$
$\prod^{n-1}_{j=0}\exp\left(\sum^{\infty}_{k=0}a_{k}(z+j)^{k}\right)=\log\left(\prod^{n-1}_{j=0}\exp\left(\sum^{\infty}_{k=0}h_{n,k}(z+j)^{k}\right)\right)\Leftrightarrow$
$\exp\left(\prod^{n-1}_{j=0}\exp\left(\sum^{\infty}_{k=0}a_{k}(z+j)^{k}\right)\right)=\prod^{n-1}_{j=0}\exp\left(\sum^{\infty}_{k=0}h_{n,k}(z+j)^{k}\right).$
Hence the problem reduces to find if for any function $X(z)$ exists function $Y_{n}(z)$ such that
$\prod^{n-1}_{j=0}X(z+j)=\sum^{n-1}_{j=0}Y_{n}(z+j)\textrm{, }\forall n\in\textbf{N}.$
However
$g(z)=\sum^{\infty}_{n=1}c_{n}\sum^{\infty}_{k=1}J_{n,k}(a_{(.)})B_{k,n}(z)=\sum^{\infty}_{n=1}\sum^{\infty}_{k=1}c_{n}J_{n,k}(a_{(.)})B_{k,n}(z).$ (a)
Assume now $e_{n}(z)$ is any base function in a set A subset of R, then
$B_{k,n}(z)=\sum^{\infty}_{l=1}S_{k,n,l}e_{l}(z)$
and relation $(a)$ becomes
$g(z)=\sum^{\infty}_{l=1}g^{*}_{l}e_{l}(z)=\sum^{\infty}_{n=1}\sum^{\infty}_{k=1}c_{n}J_{n,k}(a_{(.)})\sum^{\infty}_{l=1}S_{k,n,l}e_{l}(z)=$
$\sum^{\infty}_{l=1}\left(\sum^{\infty}_{n,k=1}c_{n}J_{n,k}(a_{(.)})S_{k,n,l}\right)e_{l}(z).$
Hence
$g^{*}_{l}=\sum^{\infty}_{n,k=1}c_{n}J_{n,k}(a_{(.)})S_{k,n,l}.$
If
$\sum^{\infty}_{n=1}I_{n,l}c_{n}=g_{l}^{*},$
where
$I_{n,l}=\sum^{\infty}_{k=1}J_{n,k}(a_{(.)})S_{k,n,l}$
and
$\sum^{\infty}_{l=1}S_{k,n,l}S^{*}_{k^{\prime},n,l}=R_{n}\delta_{k,k^{\prime}},$
then
$\sum^{\infty}_{l=1}I_{n,l}S^{*}_{k^{\prime},n,l}=\sum^{\infty}_{k=1}J_{n,k}(a_{(.)})\sum^{\infty}_{l=1}S_{k,n,l}S^{*}_{k^{\prime},n,l}=$
$\sum^{\infty}_{k=1}J_{n,k}(a_{(.)})R_{n}\delta_{k,k^{\prime}}=J_{n,k^{\prime}}(a_{(.)})R_{n}.$
Assume
$\sum^{\infty}_{n_{1},l_{1}=1}I_{nn_{1},ll_{1}}b_{n_{1},l_{1}}=\phi(n,l)$
Then
$\sum^{\infty}_{n_{1},l_{1}=1}I_{nn_{1}n_{2},ll_{1}l_{2}}b_{n_{1},l_{1}}=\phi(nn_{2},ll_{2})$
$\sum^{\infty}_{n_{1},l_{1}=1}I_{nn_{1}n_{2},ll_{1}l_{2}}b_{n_{1},l_{1}}a^{\prime}_{n_{2},l_{2}}=\phi(nn_{2},ll_{2})a^{\prime}_{n_{2},l_{2}}$
$\sum^{\infty}_{n_{2},l_{2}=1}\sum^{\infty}_{n_{1},l_{1}=1}I_{nn_{1}n_{2},ll_{1}l_{2}}b_{n_{1},l_{1}}a^{\prime}_{n_{2},l_{2}}=\sum^{\infty}_{n_{2},l_{2}=1}\phi(nn_{2},ll_{2})a^{\prime}_{n_{2},l_{2}}$
$\sum^{\infty}_{A,B=1}I_{nA,lB}\sum_{d_{1}|A,d_{2}|B}b_{d_{1},d_{2}}a^{\prime}_{A/d_{1},B/d_{2}}=\sum^{\infty}_{n_{2},l_{2}=1}\phi(nn_{2},ll_{2})a^{\prime}_{n_{2},l_{2}}$
Assume
$\sum_{d_{1}|A,d_{2}|B}b_{d_{1},d_{2}}a^{\prime}_{A/d_{1},B/d_{2}}=\delta_{A,1}\delta_{B,1}.$ (b)
Then
$I_{n,l}=\sum^{\infty}_{n_{2},l_{2}=1}\phi(nn_{2},ll_{2})a^{\prime}_{n_{2},l_{2}}$
Assume now that $b_{n_{1},l_{1}}=\delta_{l_{1},1}c^{*}_{n_{1}}$, then
$\sum^{\infty}_{n=1}\sum^{\infty}_{n_{1},l_{1}=1}I_{nn_{1},ll_{1}}b_{n_{1},l_{1}}=\sum^{\infty}_{n=1}\phi(n,l)\Rightarrow$
$\sum^{\infty}_{n,l_{1}=1}I_{n,ll_{1}}\sum_{d|n}b_{d,l_{1}}=\sum^{\infty}_{n=1}\phi(n,l)=g_{l}^{*}.$
Hence if
$\sum_{d|n}b_{d,l_{1}}=\delta_{l_{1},1}c_{n}\Leftrightarrow b_{n,l}=\delta_{l,1}\sum_{d|n}c_{d}\mu(n/d)=\delta_{l,1}c^{*}_{n},$
then
$\sum^{\infty}_{n=1}I_{n,l}c_{n}=g^{*}_{l}.$
Also if $a^{\prime}_{n,l}$ is the inverse arithmetic function of $b_{n,l}$ in the sense $(b)$, then
$\sum_{d_{1}|A,d_{2}|B}\delta_{d_{2},1}\left(\sum_{d|d_{1}}c_{d}\mu(d_{1}/d)\right)a^{\prime}_{A/d_{1},B/d_{2}}=\delta_{A,1}\delta_{B,1}\Leftrightarrow$
$\sum_{d_{1}|A}\left(\sum_{d|d_{1}}c_{d}\mu(d_{1}/d)\right)a^{\prime}_{A/d_{1},B}=\delta_{A,1}\delta_{B,1}.$ (c)
From equation $(c)$ we get the $a^{\prime}_{n,l}$. Finaly
$B_{k,n}(z)=\sum^{\infty}_{l=1}S_{k,n,l}e_{l}(z)$
$\sum^{\infty}_{l=1}S_{k,n,l}S^{*}_{k^{\prime},n,l}=R_{n}\delta_{k,k^{\prime}}$
$J_{n,k}(a_{(.)})=\frac{1}{R_{n}}\sum^{\infty}_{l=1}I_{n,l}S^{*}_{k,n,l}$
and
$I_{n,l}=\sum^{\infty}_{n_{2},l_{2}=1}\phi(nn_{2},ll_{2})a^{\prime}_{n_{2},l_{2}},$
$\sum^{\infty}_{n=1}\phi(n,l)=g^{*}_{l}.$
Assuming further that $\phi(n,l)=\phi_{1}(n)\phi_{2}(l)$, we get
$\sum^{\infty}_{n=1}\phi_{1}(n)\phi_{2}(l)=g^{*}_{l}.$
Hence
$\phi_{2}(l)=\frac{g^{*}_{l}}{G}\textrm{, }G=\sum^{\infty}_{n=1}\phi_{1}(n)$
Hence there exists constant $G$ such that
$\phi(n,l)=\phi_{1}(n)\frac{g^{*}_{l}}{G}$
and
$I_{n,l}=\frac{1}{G}\sum^{\infty}_{n_{2},l_{2}=1}\phi_{1}(nn_{2})g^{*}_{ll_{2}}a^{\prime}_{n_{2},l_{2}}$
Also then
$J_{n}(a_{k})=\frac{1}{GR_{n}}\sum^{\infty}_{l,n_{2},l_{2}=1}\phi_{1}(nn_{2})g^{*}_{ll_{2}}a^{\prime}_{n_{2},l_{2}}S^{*}_{k,n,l}$
$g(z)=\frac{1}{G}\sum^{\infty}_{n,k=1}\frac{c_{n}}{R_{n}}\sum^{\infty}_{l,n_{2},l_{2}=1}\phi_{1}(nn_{2})g^{*}_{ll_{2}}a^{\prime}_{n_{2},l_{2}}S^{*}_{k,n,l}B_{k,n}(z).$
$g(z)=\frac{1}{G}\sum^{\infty}_{n,k=1}\frac{c_{n}}{R_{n}}\sum^{\infty}_{l=1}\left(\sum^{\infty}_{n_{2},l_{2}=1}\phi_{1}(nn_{2})g^{*}_{ll_{2}}a^{\prime}_{n_{2},l_{2}}\right)S^{*}_{k,n,l}B_{k,n}(z).$
Assume that $\psi(n)$ is such that
$\sum^{\infty}_{n=1}J_{n,k}(a_{(.)})\psi(n)=\sum^{\infty}_{n=1}\frac{\psi(n)}{GR_{n}}\sum^{\infty}_{l,n_{2},l_{2}=1}\phi_{1}(nn_{2})g^{*}_{ll_{2}}a^{\prime}_{n_{2},l_{2}}S^{*}_{k,n,l}.$
Hence
$\sum^{\infty}_{n=1}J_{n,k}(a_{(.)})\psi(n)=\sum^{\infty}_{n=1}\sum^{\infty}_{l,n_{2},l_{2}=1}\frac{\psi(n)}{GR_{n}}\phi_{1}(nn_{2})g^{*}_{ll_{2}}a^{\prime}_{n_{2},l_{2}}S^{*}_{k,n,l}\Rightarrow$
$\sum^{\infty}_{n=1}J_{n,k}(a_{(.)})\psi(n)=\sum^{\infty}_{n,l=1}\phi_{1}(n)g^{*}_{l}\sum_{d_{1}|n,d_{2}|l}\frac{\psi(d_{1})}{GR_{d_{1}}}S^{*}_{k,d_{1},d_{2}}a^{\prime}_{n/d_{1},l/d_{2}}=$
$\sum^{\infty}_{n=1}\phi_{1}(n)\sum^{\infty}_{l=1}g^{*}_{l}\sum_{d_{1}|n,d_{2}|l}S^{*}_{k,d_{1},d_{2}}a^{\prime}_{n/d_{1},l/d_{2}}.$
If we set now $\phi_{1}(n)=\psi(n)=R_{n}$, ($R_{n}$ is any arithmetical function), then
$\sum^{\infty}_{n=1}J_{n,k}(a_{(.)})R_{n}=\sum^{\infty}_{n=1}R_{n}G^{-1}\sum^{\infty}_{l=1}g^{*}_{l}\sum_{d_{1}|n,d_{2}|l}S^{*}_{k,d_{1},d_{2}}a^{\prime}_{n/d_{1},l/d_{2}},$
where the $a^{\prime}_{n,m}$ is given from $(c)$. Hence
$J_{n,k}\left(a_{(.)}\right)=G^{-1}\sum^{\infty}_{l=1}g^{*}_{l}\sum_{d_{1}|n,d_{2}|l}S^{*}_{k,d_{1},b_{2}}a^{\prime}_{n/d_{1},l/d_{2}}$
Using now
$H(z)=\exp\left(\sum^{\infty}_{k=0}a_{k}z^{k}\right),$
we get
$y(x)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\prod^{\infty}_{n=0}\exp\left(-\sum^{\infty}_{k=0}J_{n,k}^{(-1)}\left(G^{-1}\sum^{\infty}_{l=1}g^{*}_{l}\sum_{d_{1}|n,d_{2}|l}S^{*}_{k,d_{1},d_{2}}a^{\prime}_{n/d_{1},l/d_{2}}\right)(z+n)^{k}\right)e^{izx}dz.$
Theorem 10.4
Given a potential
$V(z;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)e^{-i\tau nz}$
and
$B(z;\tau)=\sum^{\infty}_{n=0}b_{n}(\tau)z^{n}=g(-iz;\tau)=\sum^{\infty}_{n=0}(-i)^{n}g_{n}(\tau)z^{n},$
we form the equation
$\sum^{\infty}_{n=0}b_{n}(\tau)\frac{d^{n}}{dz^{n}}y(z;\tau)=V(z;\tau)y(z;\tau),$
which may be formed as
$V\left(i\frac{d}{dz}\right)\widehat{y}(z;\tau)=g(z;\tau)\widehat{y}(z;\tau),$
where
$\widehat{y}(z;\tau)=\int^{\infty}_{-\infty}y(x;\tau)e^{-ixz}dx.$
Also this may be written as
$\sum^{\infty}_{n=0}c_{n}(\tau)\widehat{y}(z+\tau n;\tau)=g(z;\tau)\widehat{y}(z;\tau).$ (56.1a)
I. If we assume that
$\widehat{y}(z;\tau)=\sum^{\infty}_{l=1}C_{l}(z;\tau),$ (a)
where
$C_{l}(z+n\tau;\tau)=A_{n}(z;\tau)C_{l}(z;\tau)\textrm{, }n\in\textbf{Z},$
then it must be
$g(z;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)A_{n}(z;\tau).$ (b)
From where we find $A_{n}(z;\tau)$. In this case also
$\widehat{y}(z+n\tau;\tau)=A_{n}(z;\tau)\widehat{y}(z;\tau).$ (c)
II. If $H(z;\tau)$ is such that
$\frac{\widehat{y}(z+\tau;\tau)}{\widehat{y}(z;\tau)}=H(z;\tau)\textrm{, }Im(\tau),Im(z)>0$ (d)
and $H(z+\tau;\tau)=A^{*}(z;\tau)H(z;\tau)$, then
$H(z+j\tau;\tau)=\left(\prod^{j-1}_{l=0}A^{*}(z+l\tau;\tau)\right)H(z;\tau)$
and from $(d)$:
$\widehat{y}(z+n\tau;\tau)=\left(\prod^{n-1}_{j=0}H(z+j\tau;\tau)\right)\widehat{y}(z;\tau).$ (e)
Hence from $(c)$:
$A_{n}(z;\tau)=\prod^{n-1}_{j=0}H(z+j\tau;\tau).$ (56.2)
Hence
III.
$A_{1}(z;\tau)=H(z;\tau)$
and from $(b)$ we get
$\widehat{y}(z;\tau)=\prod^{\infty}_{n=0}\frac{1}{A_{1}(z+n\tau;\tau)}.$ (f)
Proof.
For the proof of (II), we have
$g(z;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)\prod^{n-1}_{j=0}H(z+j\tau;\tau).$ (56.4)
Hence from (56.2), we get
$g(z;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)\left(\prod^{n-1}_{j=0}\prod^{j-1}_{l=0}A^{*}(z+l\tau;\tau)\right)H(z;\tau)^{n}.$
Hence given the $(-i)^{n}g_{n}(\tau)=b_{n}(\tau)$ of $B(z;\tau)$ and $c_{n}(\tau)$ of the potential, we get $H(z;\tau)$ from (56.2) and the solution of $(eq)$ is
$\widehat{y}(z;\tau)=\prod^{\infty}_{n=0}\frac{1}{H(z+n\tau;\tau)}.$
We have
$\sum^{\infty}_{k=0}b_{k}(\tau)y^{(k)}(x;\tau)=V(x;\tau)y(x;\tau)\Leftrightarrow$
$\sum^{\infty}_{k=0}b_{k}(\tau)\int^{\infty}_{-\infty}y^{(k)}(x;\tau)e^{-ix\gamma}dx=\int^{\infty}_{-\infty}y(x;\tau)\left(\sum^{\infty}_{n=0}c_{n}(\tau)x^{n}\right)e^{-ix\gamma}dx\Leftrightarrow$
$\sum^{\infty}_{k=0}b_{k}(\tau)(i\gamma)^{k}\int^{\infty}_{-\infty}y(x;\tau)e^{-ix\gamma}dx=\sum^{\infty}_{n=0}c_{n}(\tau)\int^{\infty}_{-\infty}y(x;\tau)x^{n}e^{-ix\gamma}dx\Leftrightarrow$
$B(i\gamma;\tau)\widehat{y}(\gamma)=\sum^{\infty}_{n=0}c_{n}(\tau)i^{n}\frac{d^{n}}{d\gamma^{n}}\left(\int^{\infty}_{-\infty}y(x;\tau)e^{-ix\gamma}dx\right)\Leftrightarrow$
$B(i\gamma;\tau)\widehat{y}(\gamma;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)i^{n}\frac{d^{n}}{d\gamma^{n}}\widehat{y}(\gamma;\tau)\Leftrightarrow$
$B(i\gamma;\tau)\widehat{y}(\gamma;\tau)=V\left(i\frac{d}{d\gamma}\right)\widehat{y}(\gamma;\tau).$ (57)
Assume now that $V(x;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)e^{-in\tau x}$. Hence
$B(i\gamma;\tau)\widehat{y}(\gamma;\tau)=\sum^{\infty}_{n=0}c_{n}(\tau)\exp\left(\tau n\frac{d}{d\gamma}\right)\widehat{y}(\gamma;\tau).$ (58)
Or equivalent
$B(i\gamma;\tau)\widehat{y}(\gamma;\tau)=\sum^{\infty}_{n=-\infty}c_{n}(\tau)\widehat{y}(\gamma+n\tau;\tau).$ (59)
Assume that $Im(z)$, $Im(\tau)>0$ and
$\widehat{y}(z;\tau)=\sum^{\infty}_{l=1}C_{l}(z;\tau).$ (60)
Then if $n$ is integer, we have
$C_{l}(z+n\tau;\tau)=A_{n}(z;\tau)C_{l}(z;\tau).$
Hence for almost all $a_{n}(\tau)$ we have
$\sum^{\infty}_{n=0}a_{n}(\tau)\widehat{y}(z+n\tau)=\sum^{\infty}_{n=0}a_{n}(\tau)\sum^{\infty}_{l=1}C_{l}(z+n\tau;\tau)=$
$\sum^{\infty}_{n=0}\sum^{\infty}_{l=1}a_{n}(\tau)A_{n}(z;\tau)C_{l}(z;\tau)=\sum^{\infty}_{n=0}a_{n}(\tau)A_{n}(z;\tau)\sum^{\infty}_{l=1}C_{l}(z;\tau)=$
$\left(\sum^{\infty}_{n=0}a_{n}(\tau)A_{n}(z;\tau)\right)\widehat{y}(z;\tau).$
Hence we can write
$\sum^{\infty}_{n=0}a_{n}(\tau)\left(\widehat{y}(z+n\tau;\tau)-A_{n}(z;\tau)\widehat{y}(z;\tau)\right)=0$
and it must be
$\widehat{y}(z+n\tau;\tau)=A_{n}(z;\tau)\widehat{y}(z;\tau).$ (61)
Hence if $a_{n}(\tau)=c_{n}(\tau)$ and
$g(z;\tau)=B(iz;\tau)=\sum^{\infty}_{n=0}b_{n}(\tau)(iz)^{n}=\sum^{\infty}_{n=0}c_{n}(\tau)A_{n}(z;\tau),$
then from (d) $\widehat{y}(z)$ is given from (f) and satisfies (56.1a) and $y(x;\tau)$ is solution of
$\sum^{\infty}_{n=0}b_{n}(\tau)y^{(n)}(x;\tau)=V(x;\tau)y(x;\tau),$ (62)
where
$V(x)=\sum^{\infty}_{n=0}c_{n}(\tau)e^{-in\tau x}.$
5 A relativistic wave equation
We have
$\left[\widehat{E},\widehat{U}\right]\Psi=\left(\widehat{E}\widehat{U}\Psi-\widehat{U}\widehat{E}\Psi\right)=i\hbar\partial_{t}\left(\widehat{U}\Psi\right)-i\hbar\widehat{U}\partial_{t}\Psi=i\hbar\left(\partial_{t}\widehat{U}\right)\Psi.$
Hence
Lemma 1.
It holds in general (for any operator $\widehat{U}$ and any function $\Psi$):
$i\hbar\left(\partial_{t}\widehat{U}\right)\Psi=\left[\widehat{E},\widehat{U}\right]\Psi.$
Theorem 11.
Assume the equation
$E\Psi=H\Psi$ (eq)
Assume operator $U$ satisfies the equation
$i\hbar\left(\partial_{t}U\right)=\left[H,U\right],$ (eq1)
where $H$ is the self adjoint Hamiltonian of $(eq)$. Then
I. $U$ sends a solution $\Psi$ of $(eq)$ to $U\Psi$, which is also a solution of $(eq)$.
II. $U$ is almost self adjoint i.e. for every $\Psi,\Phi$ solutions of $(eq)$ exists constant $c$ such that:
$\left(U\Phi,\Psi\right)-\left(\Phi,U\Psi\right)=c=\textrm{constant}.$
III. The operator $U$ is almost self-adjoint operator and also invariant i.e. is a conserved quantity of the system.
Remarks. Equation $(eq1)$ is equivalent to $(\textbf{I})$.
Proof.
It holds for every operator $U$ the relation
$i\hbar\left(\partial_{t}U\right)=\left[E,U\right].$
Hence from $(eq1)$, we get easily that if $\Psi$ is solution of $E\Psi=H\Psi$, then
$\left[U,H\right]\Psi+\left[E,U\right]\Psi=0\Leftrightarrow$
$UH\Psi-HU\Psi+EU\Psi-UE\Psi=0\Leftrightarrow$
$EU\Psi=HU\Psi.$
Hence $\Psi^{\prime}=U\Psi$ is clearly solution of $(eq)$. The inverse of this is also easily proved.
If $\Phi$ and $\Psi$ are solutions of $(eq)$, then differentiating $(U\Phi,\Psi)-(\Phi,U\Psi)$ with respect to $t$, we arive equivalently to the equation
$\partial_{t}\left(\left(U\Phi,\Psi\right)-\left(\Phi,U\Psi\right)\right)=0.$
Hence easily we conclude that
$\left(U\Phi,\Psi\right)-\left(\Phi,U\Psi\right)=c=\textrm{constant.}$
$U$ is almost self adjoint.
It is also easy to see someone that $i\hbar\frac{d}{dt}\left\langle U\right\rangle=0$. Hence $U$ is a conserved quantity.
Examples.
Assume that the Hamiltonian is independent of time $t$ (for example $H=\left(p\right)^{2}/(2m)+V(x)$). Then from
$i\hbar\partial_{t}H=0=\left[H,H\right]=0,$
we get that if $\Psi$ is solution of $(eq)$, then $H\Psi$ is also a solution of $(eq)$. Also then
$\left\langle E\right\rangle=\left\langle H\right\rangle=\textbf{constant.}$
Hence the energy is conserved. The same holds if we replace
$U=f(H)=\sum^{\infty}_{k=0}a_{k}H^{k}.$
Then if $\Psi$ is solution of $(eq)$, the quantity $f(H)\Psi$ is also solution of $(eq)$. Also then $f(H)$ is almost self adjoint and conserved quantity. Note that
$i\hbar\left(\partial_{t}f(H)\right)=0=\left[H,f(H)\right].$
Assume that we have two equations
$i\hbar\frac{\partial}{\partial t}\Psi(x,t)=H\Psi(x,t)\textrm{ and }i\hbar\frac{\partial}{\partial x}\Psi(x,t)=W\Psi(x,t).$ (55)
Here $x,t$ are the variables and $\Psi=\Psi(x,t)$ is the solution of our system. We now introduce a Lorentz boost along $x$:
$x=\gamma_{0}x^{\prime}+\gamma_{0}vt^{\prime}\textrm{ and }t=\gamma_{0}vc^{-2}x^{\prime}+\gamma_{0}t^{\prime},$ (56)
where $\gamma_{0}=(1-v^{2}/c^{2})^{-1/2}$. The inverse of this transformation is
$x^{\prime}=\gamma_{0}x-\gamma_{0}vt\textrm{ and }t^{\prime}=-\gamma_{0}vc^{-2}x+\gamma_{0}t$ (57)
The derivatives of all functions $F(x,t)$ are obeying
$\frac{\partial}{\partial t}=-\gamma_{0}v\frac{\partial}{\partial x^{\prime}}+\gamma_{0}\frac{\partial}{\partial t^{\prime}}$ (58)
and
$\frac{\partial}{\partial x}=\gamma_{0}\frac{\partial}{\partial x^{\prime}}-\gamma_{0}vc^{-2}\frac{\partial}{\partial t^{\prime}}$ (59)
and the inverse is
$\frac{\partial}{\partial t^{\prime}}=\gamma_{0}v\frac{\partial}{\partial x}+\gamma_{0}\frac{\partial}{\partial t}$ (60)
and
$\frac{\partial}{\partial x^{\prime}}=\gamma_{0}\frac{\partial}{\partial x}+\gamma_{0}vc^{-2}\frac{\partial}{\partial t}.$ (61)
Before the Lorentz transformation our equations where that of (57). After, they become
$i\hbar\frac{\partial}{\partial t^{\prime}}\Psi^{\prime}(x^{\prime},t^{\prime})=H^{\prime}\Psi^{\prime}(x^{\prime},t^{\prime})\textrm{ and }i\hbar\frac{\partial}{\partial x^{\prime}}\Psi^{\prime}(x^{\prime},t^{\prime})=W^{\prime}\Psi^{\prime}(x^{\prime},t^{\prime}).$ (62)
We define the operator $S$ which sends $\Psi(x,t)$ to $\Psi^{\prime}(x^{\prime},t^{\prime})$, such:
$\Psi^{\prime}(x^{\prime},t^{\prime})=S\Psi(x,t).$ (63)
Then
$i\hbar\frac{\partial}{\partial t}\Psi=H\Psi\Leftrightarrow i\hbar\frac{\partial}{\partial t}\left(S^{-1}\Psi^{\prime}\right)=H\left(S^{-1}\Psi^{\prime}\right).$ (64)
Hence
$I=i\hbar\frac{\partial}{\partial t}\Psi=i\hbar\frac{\partial}{\partial t}\left(S^{-1}\Psi^{\prime}\right)=i\hbar\left(\frac{\partial}{\partial t}S^{-1}\right)\Psi^{\prime}+S^{-1}i\hbar\frac{\partial}{\partial t}\Psi^{\prime}=$
$i\hbar\gamma_{0}\left(\frac{\partial S^{-1}}{\partial t^{\prime}}\right)\Psi^{\prime}-\gamma_{0}vi\hbar\left(\frac{\partial S^{-1}}{\partial x^{\prime}}\right)\Psi^{\prime}+\gamma_{0}i\hbar S^{-1}\frac{\partial\Psi^{\prime}}{\partial t^{\prime}}-i\hbar\gamma_{0}vS^{-1}\frac{\partial\Psi^{\prime}}{\partial x^{\prime}}=$
$\gamma_{0}\left(i\hbar\left(\frac{\partial S^{-1}}{\partial t^{\prime}}\right)+S^{-1}H^{\prime}\right)\Psi^{\prime}-\gamma_{0}v\left(i\hbar\left(\frac{\partial S^{-1}}{\partial x^{\prime}}\right)+S^{-1}W^{\prime}\right)\Psi^{\prime}.$
At this point we define the derivatives (operators)
$\left(D_{t}P\right)\Phi=i\hbar\left(\frac{\partial P}{\partial t}\right)\Phi+P(H\Phi)$ (65)
and
$\left(D_{x}P\right)\Phi=i\hbar\left(\frac{\partial P}{\partial x}\right)\Phi+P(W\Phi),$ (66)
for all operatrors $P$ and functions $\Phi$. To see someone this more simply we have
$\left(D_{t}P\right)\Phi=E(P\Phi)=i\hbar\frac{\partial}{\partial t}\left(P\Phi\right)=(EP)\Phi+P(E\Phi)=$
$i\hbar\left(\frac{\partial P}{\partial t}\right)\Phi+P\left(E\Phi\right).$ (67)
In the case of $\Phi=\Psi$ is the solution of $E\Psi=H\Psi$, we have
$\left(D_{t}P\right)\Psi=i\hbar\left(\frac{\partial P}{\partial t}\right)\Psi+P(H\Psi).$ (68)
In the same way
$\left(D_{x}P\right)\Psi=i\hbar\left(\frac{\partial P}{\partial x}\right)\Psi+P(W\Psi).$ (69)
Also $D_{t}$ sends an operator $P$ to operator $D_{t}P$:
$D_{t}(.)=i\hbar\frac{\partial(.)}{\partial t}+(.)H$ (70)
and
$D_{x}(.)=i\hbar\frac{\partial(.)}{\partial x}+(.)W.$ (71)
We can also set
$D_{t^{\prime}}(.)=i\hbar\frac{\partial(.)}{\partial t^{\prime}}+(.)H^{\prime}$ (72)
and
$D_{x^{\prime}}(.)=i\hbar\frac{\partial(.)}{\partial x^{\prime}}+(.)W^{\prime}.$ (73)
We can assume $D_{t}P$ and $D_{x}P$ have the properties of (60) and (61) i.e. we have
$D_{t}=\gamma_{0}D_{t^{\prime}}-\gamma_{0}vD_{x^{\prime}}\textrm{ and }D_{x}=-\gamma_{0}vc^{-2}D_{t^{\prime}}+\gamma_{0}D_{x^{\prime}}.$ (74)
This happens if and only if the operators $H^{\prime},W^{\prime}$ are
$H^{\prime}=\gamma_{0}H+v\gamma_{0}W\textrm{ and }W^{\prime}=v\gamma_{0}c^{-2}H+\gamma_{0}W.$ (75)
Hence $I$ becomes
$I=\gamma_{0}\left(D_{t^{\prime}}S^{-1}\right)\Psi^{\prime}-\gamma_{0}v\left(D_{x^{\prime}}S^{-1}\right)\Psi^{\prime}=\left(D_{t}S^{-1}\right)\Psi^{\prime}=$
$i\hbar\frac{\partial S^{-1}}{\partial t}\Psi^{\prime}+S^{-1}H\Psi^{\prime}.$
Hence from (63):
$i\hbar\frac{\partial S^{-1}}{\partial t}\Psi^{\prime}+S^{-1}\left(H\Psi^{\prime}\right)=H\left(S^{-1}\Psi^{\prime}\right)\Leftrightarrow$
$i\hbar\frac{\partial S^{-1}}{\partial t}\Psi^{\prime}=\left[H,S^{-1}\right]\Psi^{\prime}.$ (76)
Also in the same way
$i\hbar\frac{\partial S^{-1}}{\partial x}\Psi^{\prime}=\left[W,S^{-1}\right]\Psi^{\prime}.$ (77)
This is the case of Theorem 11, with $U=S^{-1}$. However $\Psi^{\prime}=S\Psi$.
Hence we get the next
Theorem 12.
Assume that we have two equations
$i\hbar\frac{\partial}{\partial t}\Psi(x,t)=H\Psi(x,t)\textrm{ and }i\hbar\frac{\partial}{\partial x}\Psi=W\Psi(x,t).$ (78)
Here $x,t$ are the variables and $\Psi=\Psi(x,t)$ is the solution of our system. We now introduce a Lorentz boost along $x$:
$x=\gamma_{0}x^{\prime}+\gamma_{0}vt^{\prime}\textrm{ and }t=\gamma_{0}vc^{-2}x^{\prime}+\gamma_{0}t^{\prime},$ (79)
where $\gamma_{0}=(1-v^{2}/c^{2})^{-1/2}$. Before the Lorentz transformation our equations where that of (80). After, they become
$i\hbar\frac{\partial}{\partial t^{\prime}}\Psi^{\prime}(x^{\prime},t^{\prime})=H^{\prime}\Psi^{\prime}(x^{\prime},t^{\prime})\textrm{ and }i\hbar\frac{\partial}{\partial x^{\prime}}\Psi^{\prime}(x^{\prime},t^{\prime})=W^{\prime}\Psi^{\prime}(x^{\prime},t^{\prime}).$ (80)
We define the operator $S$ which sends $\Psi(x,t)$ to $\Psi^{\prime}(x^{\prime},t^{\prime})$, such that:
$\Psi^{\prime}(x^{\prime},t^{\prime})=S\Psi(x,t).$ (81)
Hence from (78),(79), we have
$E\Psi=H\Psi\textrm{ and }-p\Psi=W\Psi$ (82)
and
$S^{-1}\left(E-H\right)S\Psi=0\textrm{ and }S^{-1}\left(p+W\right)S\Psi=0.$ (83)
Continuing we can write
$E^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{x^{\prime}}=i\hbar\frac{\partial}{\partial t^{\prime}}\int^{\infty}_{-\infty}\overline{\Psi^{\prime}}\Psi^{\prime}dx^{\prime}=$
$i\hbar\frac{\partial}{\partial t^{\prime}}\left(\int^{\infty}_{-\infty}\overline{\Psi^{\prime}}\Psi^{\prime}(-\gamma_{0}v)dt+\int^{\infty}_{-\infty}\overline{\Psi^{\prime}}\Psi^{\prime}\gamma_{0}dx\right)=$
$(\gamma_{0}v)^{2}i\hbar\frac{\partial}{\partial x}\int^{\infty}_{-\infty}\overline{\Psi^{\prime}}\Psi^{\prime}dt+\gamma_{0}^{2}i\hbar\frac{\partial}{\partial t}\int^{\infty}_{-\infty}\overline{\Psi^{\prime}}\Psi^{\prime}dx.$
Hence
$E^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{x^{\prime}}=-(\gamma_{0}v)^{2}\left(\left(p\Psi^{\prime},\Psi^{\prime}\right)_{t}-\left(\Psi^{\prime},p\Psi^{\prime}\right)_{t}\right)$-
$\gamma_{0}^{2}\left(\left(E\Psi^{\prime},\Psi^{\prime}\right)_{x}-\left(\Psi^{\prime},E\Psi^{\prime}\right)_{x}\right).$ (84)
However if $H,W$ are self-adjoint, then from Theorem 12:
$\left(p\Psi^{\prime},\Psi^{\prime}\right)_{t}-\left(\Psi^{\prime},p\Psi^{\prime}\right)_{t}=\int_{R}\overline{(pS\Psi)}S\Psi dt-\int_{R}\overline{S\Psi}p(S\Psi)dt=$
$\int_{R}\overline{S^{+}p(S\Psi)}\Psi dt-\int_{R}\overline{\Psi}S^{+}(p(S\Psi))dt=$
$\int_{R}\overline{S^{+}WS\Psi}\Psi dt+\int_{R}\overline{\Psi}S^{+}W(S\Psi)dt=$
$\int_{R}\overline{\Psi}S^{+}WS\Psi dt+\int_{R}\overline{\Psi}S^{+}WS\Psi dt=0.$
In the same way we find
$\left(E\Psi^{\prime},\Psi^{\prime}\right)_{x}-(\Psi^{\prime},E\Psi^{\prime})_{x}=0,$
from the self-adjointnes of $H$ and Theorem 12. Hence (if $H,W$ are self adjoint), then
$E^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{x^{\prime}}=0$ (85)
Also
$p^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{t^{\prime}}=-(\gamma_{0}vc^{-2})^{2}\left(\left(E\Psi^{\prime},\Psi^{\prime}\right)_{x}-\left(\Psi^{\prime},E\Psi^{\prime}\right)_{x}\right)+$
$\gamma_{0}^{2}\left(\left(p\Psi^{\prime},\Psi^{\prime}\right)_{t}-\left(\Psi^{\prime},p\Psi^{\prime}\right)_{t}\right).$ (86)
Hence when $H,W$ are self-adjoint, then
$p^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{t^{\prime}}=0.$ (87)
Proposition.
If the function $\Psi(x,t)$ satisfies (84) and (85), with $H,W$ self-adjoint, then
$E^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{x^{\prime}}=p^{\prime}\left(\Psi^{\prime},\Psi^{\prime}\right)^{\prime}_{t^{\prime}}=0,$ (88)
where
$\Psi^{\prime}(x^{\prime},t^{\prime})=S\Psi(x,t).$ (89)
Assume now the equations $E\Psi=H\Psi:(eq_{1})$ and $p\Psi=W\Psi:(eq_{2})$. Then making a Lorentz transformation $\Lambda$ in $(eq_{1}),(eq_{2})$, we have $E^{\prime}\Psi^{\prime}=H^{\prime}\Psi^{\prime}:(eq^{\prime}_{1})$, $p^{\prime}\Psi^{\prime}=W^{\prime}\Psi^{\prime}:(eq^{\prime}_{2})$ and $\Psi^{\prime}=S\Psi$. We know from above evaluations, that if $\Psi$ is solution of $(eq_{1}),(eq_{2})$ and $\Psi^{\prime}$ is a solution of $(eq^{\prime}_{1}),(eq^{\prime}_{2})$, then $\left(\Psi,\Psi\right)_{x}=\left(\Psi,\Psi\right)_{t}=1$ and $\left(\Psi^{\prime},\Psi^{\prime}\right)_{x}=\left(\Psi^{\prime},\Psi^{\prime}\right)_{t}=1$. Hence
$\left(\Psi^{\prime},\Psi^{\prime}\right)_{x}=c_{3}\left(\Psi,\Psi\right)_{x}=c_{3}\Leftrightarrow\left(S\Psi,S\Psi\right)_{x}=c_{3}\left(\Psi,\Psi\right)_{x}\Leftrightarrow$
$\left(\Psi,S^{+}S\Psi\right)_{x}=(\Psi,c_{3}\Psi)_{x}\Leftrightarrow\left(\Psi,\{S^{+}S-c_{3}I\}\Psi\right)_{x}=0,$
for all $\Psi\in L$. Also
$\left(\Psi,\{S^{+}S-c_{4}I\}\Psi\right)_{t}=0.$
All these, are in the sense that an operator $S$ is unitary iff (in a inner product) holds
$\left(S\Psi,S\Phi\right)_{x}=\left(\Psi,\Phi\right)_{x}\textrm{, }\forall\Psi,\Phi\in M.$ (90)
and
$\left(S\Psi,S\Phi\right)_{t}=\left(\Psi,\Phi\right)_{t}\textrm{, }\forall\Psi,\Phi\in M.$ (91)
Hence if $\Psi^{\prime}=S\Psi$, $\Psi=S^{-1}\Psi^{\prime}$, $S^{+}S\Psi=\Psi$, we have:
Theorem 13.
Assume the general equations $E\Psi=H\Psi:(eq_{1})$, $p\Psi=W\Psi:(eq_{2})$. Making in $(eq_{1}),(eq_{2})$, a Lorentz transformation of variables, we get $E^{\prime}\Psi^{\prime}=H^{\prime}\Psi^{\prime}:(eq^{\prime}_{1})$, $p^{\prime}\Psi^{\prime}=W^{\prime}\Psi^{\prime}:(eq^{\prime}_{2})$. Assuming the operator $S$ is such that $\Psi^{\prime}=S\Psi$, then if $H,W$ of Theorem 12, are self-adjoint, we can normalize $\Psi,\Psi^{\prime}$, and write $(\Psi^{\prime},\Psi^{\prime})_{x}=(\Psi,\Psi)_{x}=1$ and $\left(\Psi^{\prime},\Psi^{\prime}\right)_{t}=\left(\Psi,\Psi\right)_{t}=1$. Hence
$S^{+}\Psi=S^{-1}\Psi\textrm{, }\forall\Psi\in L.$ (92)
Theorem 14.
A relativistic wave system obey the following equations:
$E\Psi=H\Psi\textrm{ and }-p\Psi=W\Psi$ (93)
$i\hbar\frac{\partial S^{-1}}{\partial t}\Psi^{\prime}=\left[H,S^{-1}\right]\Psi^{\prime}$ (94)
and
$i\hbar\frac{\partial S^{-1}}{\partial x}\Psi^{\prime}=\left[W,S^{-1}\right]\Psi^{\prime}.$ (95)
The operator $S$ is unitary (that of Theorems 12,13). The functions $\Psi=\Psi(x,t)$ which are solutions of (84),(85) (and equivalent to (96),(97)), form the set of solutions $L$. The more general space $M$, with two inner products
$\left(f,g\right)_{x}:=\int^{\infty}_{-\infty}\overline{f(x,t)}g(x,t)dx$ (96)
and
$\left(f,g\right)_{t}:=\int^{\infty}_{-\infty}\overline{f(x,t)}g(x,t)dt,$ (97)
is considered as the Hilbert space of our system. The space $L$ also have two inner products as being subspace of $M$.